### v1

# ConsDreamer: Advancing Multi-View Consistency for Zero-Shot Text-to-3D Generation

Abstract
Recent advances in zero-shot text-to-3D generation have revolutionized 3D content creation by enabling direct synthesis from textual descriptions. While state-of-the-art methods leverage 3D Gaussian Splatting with score distillation to enhance multi-view rendering through pre-trained text-to-image (T2I) models, they suffer from inherent view biases in T2I priors. These biases lead to inconsistent 3D generation, particularly manifesting as the multi-face Janus problem, where objects exhibit conflicting features across views. To address this fundamental challenge, we propose ConsDreamer, a novel framework that mitigates view bias by refining both the conditional and unconditional terms in the score distillation process: (1) a View Disentanglement Module (VDM) that eliminates viewpoint biases in conditional prompts by decoupling irrelevant view components and injecting precise camera parameters; and (2) a similarity-based partial order loss that enforces geometric consistency in the unconditional term by aligning cosine similarities with azimuth relationships. Extensive experiments demonstrate that ConsDreamer effectively mitigates the multi-face Janus problem in text-to-3D generation, outperforming existing methods in both visual quality and consistency.
Index Terms:
text to 3D generation, multi-face Janus problem, score distillation sampling.
publicationid: pubid: Copyright ©20xx IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.
I Introduction
[…]
[lines=2]3D GENERATION technology plays a crucial role in various fields such as innovative industrial design, game development, and virtual reality. In particular, zero-shot text-to-3D generation [1], [2], [3], [4], [5] aims to generate 3D content without 3D training data, enabling the conversion from concept to reality. However, zero-shot text-to-3D generation tasks [6], [7], [8], [9] are constrained by the inherent complexity of the wild world and the scarcity of 3D data, unlike text-to-image (T2I) tasks [10], [11]. From this perspective, generating high-quality 3D content from text is still a significant challenge.
Earlier works [12], [13], [14], [15], [16] based on Neural Radiance Fields (NeRF) introduce a framework that extended the 2D generative capabilities of diffusion models into the 3D domain by leveraging a view-conditioned projection mechanism. For instance, DreamFusion [14] introduces the Score Distillation Sampling (SDS) technique, effectively aligning the rendering of 3D content with 2D generative priors. However, rendering in NeRF primarily relies on implicit 3D representations, which significantly increases computational cost and makes it challenging to produce high-resolution images [17]. Recently, 3D Gaussian Splatting [18] has been introduced as an efficient alternative to NeRF’s implicit 3D representation, enabling real-time, high-resolution rendering through explicit point-based modelling. Its adoption in SOTA zero-shot text-to-3D generation methods [17], [19], [20], [21] highlights its effectiveness in delivering superior results in quality and speed. However, these recent methods still struggle with generation precision, e.g., the multi-face Janus problem [13], [14]. A single 3D object exhibiting inconsistencies across its faces in different views, as shown in Fig. 1, significantly undermines the realism and coherence of the 3D output. Therefore, ensuring view consistency in 3D generation is crucial for high-quality results, which is the primary motivation for this paper.
Figure 1: Examples of Multi-Face Janus Problem: (a) generated by DreamFusion [14], (b) by Sherpa3D [22], (c) by LucidDreamer [17], (d) by DreamScene [23], and (e) by ScaleDreamer [24].
Figure 2: Examples of text-to-3D content creation using our framework. We present a text-to-3D generation framework, named $ConsDreamer$, which leverages a View Disentanglement Model and a novel partial order loss to ensure semantic clarity across views (detailed in Section III). The generative 3D results demonstrate the superiority of ConsDreamer. Please zoom in for details.
Considering the most recent text-to-3D methods, they predominantly rely on T2I models to generate reference images. However, these models are typically trained on datasets sourced from publicly available resources and online platforms, such as photo-sharing sites. A notable limitation of these datasets is the lack of diverse and well-annotated multi-view images, which results in a bias toward high-frequency single view text-image pairs. As a consequence, T2I models tend to prioritise certain viewpoints, particularly frontal or side views [25]. This view bias adversely impacts the consistency of 3D generation. When given rendered images and corresponding textual descriptions specifying particular views, the model often defaults to its prior view preference, rather than adhering to the exact view that training required. This misalignment leads to multi-view inconsistencies, where features across different views are incompatible, exacerbating the multi-face Janus problem in text-to-3D generation. Furthermore, the dominance of preferred viewpoints introduces an unbalanced view distribution, where even in the absence of explicit view conditions, the model still favours certain perspectives.
To tackle the aforementioned problem, we first conducted a mathematical analysis of the Janus problem. The analysis reveals that view biases arise during parameter updates, affecting both conditional and unconditional terms used in existing text-to-3D methods. In conditional terms, the input prompts guide view control, but conflicts between prior view preferences and target view specifications result in ambiguous view semantics. Unconditional terms, independent of prompts, are directly shaped by inherent view biases. Therefore, this paper proposes a novel framework, ConsDreamer, to address the challenges from both conditional and unconditional perspectives. For the conditional term, we propose the View Disentanglement Module (VDM), which starts extracting view-agnostic keyword embeddings from the prompt by combining the extracted keyword with explicit view descriptions (e.g., “front view,” “side view”). These view-specific embeddings are processed to isolate view-related features by subtracting their projection from the view-agnostic embedding space. The resulting view-specific features are used to eliminate prior view preferences and inject target view information, ensuring that the generated content adheres to user-specified views.
For the unconditional term, we introduce a similarity-based partial order loss $\mathcal{L}_{P}$, which leverages the inherent relationship between view similarity and azimuthal angle distances. To implement this, a Cartesian coordinate framework is established, and an expected similarity partial order is determined based on the proximity of the azimuthal angles of multiple views to a reference view. The actual cosine similarity scores of the rendered images are constrained to conform to this partial order. By enforcing these constraints, $\mathcal{L}_{P}$ ensures cross-view consistency in 3D content, effectively addressing the multi-face Janus problem. We also conducted comprehensive experiments to illustrate the efficacy of our method, confirming that ConsDreamer can better alleviate the multi-face Janus problem and strengthen semantic correlations across different views and enhance generated content quality. The main contributions of this work are summarized as follows:
  • $\bullet$
    We begin by examining the fundamental principles of T2I models to formulate a mathematical framework for text-to-3D generation. This framework provides an in-depth analysis of the root causes of the multi-face Janus problem. Our analysis identifies two key problematic terms—conditional and unconditional—that introduce prior view biases during the score distillation process. To address these issues, we propose optimization strategies for both terms.
  • $\bullet$
    We propose a View Disentanglement Model (VDM) to eliminate prior view preferences and integrate target view control in the conditional term. For the unconditional term, we introduce a similarity-based partial order loss to enhance the model’s view-awareness. Together, these components synergistically improve the clarity of view semantics.
  • $\bullet$
    We conduct extensive experiments on a wide range of prompts to validate the effectiveness of the proposed method.
II Related Works
II-A Differentiable 3D Representations
Differentiable 3D representation methods are especially essential for zero-shot text-guided 3D generation. Their differentiable nature enables a trainable pathway, optimizing 3D parameters by ensuring that rendered results from random views align with the prompt inputs. Previously, various representations are introduced for text-to-3D generation [12], [26], [27], [28]. Among these, NeRF [12] emerges as the most widely adopted representation in text-to-3D generation tasks. NeRF enables novel view synthesis from unobserved viewpoints and leverages guidance information from the T2I model to optimize the tri-plane [29] parameters through a pre-trained MLP network, demonstrating broad applicability across downstream 3D tasks [3], [12], [13], [14], [15]. Despite its success, NeRF faces challenges in rendering speed and memory consumption due to its implicit nature. A new 3D representation method, named 3D Gaussian Splatting, is thus introduced. Gaussian Splatting-based methods [18], [17], [30], [22] address these limitations using explicit 3D representations with anisotropic Gaussians. They ensure high reconstruction quality, real-time rendering, and reduced computational demands, making them ideal for large-scale and dynamic scenes. By replacing the NeRF method with the explicit 3D Gaussian Splatting representation, it becomes possible to integrate pre-trained 3D point cloud models like Point-E [31] and Shape-E [32] into the 3D generation framework, providing a prior 3D knowledge foundation.
II-B Diffusion models
Diffusion models serve as a core component for guiding text-to-3D generation. The Denoising Diffusion Probabilistic Model (DDPM) [33], based on Markov chains, operates by gradually “diffusing” data from its original state into random noise in the forward procedure and then reversing the process to reconstruct the original data. Although DDPM achieves “diffusion-reconstruction,” its efficiency is hindered by the step-by-step sampling strategy. Subsequently, Denoising Diffusion Implicit Models (DDIM) [11] introduce a deterministic sampling approach, enabling high-quality generation with significantly fewer sampling steps. Additionally, DDPM faces high time consumption as the data retains its original size at every step. Latent Diffusion Models (LDM) [34] address this limitation by transferring the diffusion process from the original pixel space to a low-dimensional latent space, significantly improving efficiency, where the attention [7], [35] mechanism could help to control the learning process. Most recently, Stable Diffusion [34] introduces a conditional generation approach based on the LDM, enabling high-quality image generation from textual descriptions. Subsequent research [17], [36], [37] extends Stable Diffusion with the deterministic sampling mechanism in DDIM.
II-C Text-to-3D Generation
Early efforts, such as DreamField [38], leverage the cross-modal information of CLIP [39] to translate text into 3D. However, DreamField faces limitations due to the lack of high-precision alignment in CLIP’s cross-modal information, resulting in suboptimal 3D generation performance. With the advent of diffusion models, a new paradigm [14], [30] emerges for extracting 3D assets from pre-trained T2I models, leveraging 2D diffusion model outputs to guide differentiable 3D representations. For example, DreamFusion introduces Score Distillation Sampling (SDS), which aligns 3D rendering information with diffusion priors and optimizes model parameters through backpropagation. However, SDS-based approaches often introduce smoothing effects. Recent research [40], [41], [42], [43] aims to address this. Prolific-Dreamer [41] proposes Variational Score Distillation (VSD), significantly improving generation quality. Consistent 3D [9] analyzes SDS from ordinary differential equations (ODEs) and introduces Consistency Distillation Sampling (CSD). ScaleDreamer [24] introduces Asynchronous Score Distillation (ASD), which shifts diffusion timesteps to minimize noise prediction errors without fine-tuning the diffusion model. LucidDreamer [17] addresses over-smoothing issues via Interval Score Matching (ISM), integrating DDIM inversion to enable deterministic sampling. Subsequent advancements, DreamerXL [37], ExactDreamer [36], and Guided Consistency Sampling (GCS) [21] further refine this paradigm by systematically reducing accumulated errors in DDIM’s reverse process. The methods above are predominantly optimization-based, while recent advances also explore feed-forward methodologies [44], [45], [46] to facilitate the generation of large-scale 3D assets efficiently [47] or to integrate auxiliary information such as text-to-3D knowledge graphs [48]. These methods, however, demand substantial GPU resources for additional training and require extensive 3D datasets.
Figure 3: An overview of ConsDreamer. Our framework is built upon the main flow of 3D content distilled from the T2I model. ConsDreamer introduces two key innovations: (a) VDM disentangles the keyword in the prompt to obtain the canonical view features $\delta_{\text{key}}^{\text{view}}$, which are then used for precise view control (detailed in Section III-B). (b) A novel partial order loss $\mathcal{L}_{P}$ is introduced among multi-view rendered images to endow the model with view-aware capabilities (detailed in Section III-C). Together, the VDM and $\mathcal{L}_{P}$ enhance the clarity of view semantics and significantly mitigate the multi-face Janus problem.
Several works specifically address the multi-face Janus problem. Perp-Neg [25] leverages the geometric properties of score space to enhance the utilization of negative prompts. MVDream [13] fine-tunes pre-trained T2I models using canonical view image combinations, ensuring view consistency but requiring additional training and datasets. Debiased-SDS [15] mitigates the multi-face Janus problem in two ways: first, by clipping distillation scores to suppress erroneous artifacts, and second, by identifying and omitting words in the prompt that conflict with the target view. However, Debiased-SDS lacks fine-grained view control, and omitting conflicting words can compromise the integrity of user prompts. Sherpa3D [22] uses structural alignment and semantic guidance, incorporating normal map gradients and CLIP semantic embeddings to preserve the geometric consistency of the initial 3D prior during 2D lifting optimization. However, this method heavily relies on coarse 3D prior information. Overall, existing approaches lack precise understanding and control of view-related information while requiring additional computational resources. In this paper, we introduce a novel framework that eliminates prior view biases by disentangling view features and ensuring multi-view consistency through a carefully designed partial order loss, significantly enhancing view semantic clarity without requiring additional training.
III Methodology
To tackle the multi-face Janus problem, we propose ConsDreamer, a unified framework to enhance view semantic clarity. Fig.3 provides an overview of our approach. Section III-A presents a mathematical analysis of the problem, identifying two problematic terms: the conditional and unconditional terms, both of which introduce prior view preferences during the score distillation process. Based on this analysis, Section III-B introduces VDM to eliminate prior view preferences and integrate target view control in the conditional term. Meanwhile, for the unconditional term, Section III-C explores similarity distribution patterns across views in rendered images and proposes a similarity-based partial order loss $\mathcal{L}_{P}$ to enhance spatial consistency in 3D content.
III-A Analysis of the multi-face Janus problem
In the Stable Diffusion framework, image generation is formulated as a reverse process that reconstructs the initial state $z_{0}$ from the final noise $z_{T}$. Guided by user prompts $c$, this reverse recovery models the probability distribution $p_{2D}(z_{0}\mid c)$ for generating the final 2D latent representation $z_{0}$, as expressed in the following equation:
$p_{2D}\left(\boldsymbol{z}_{0}\!\mid\!c\right)\!=\!\int\!p\left(\boldsymbol{z}_{T}\right)\!\prod_{t=1}^{T}\!p\left(\boldsymbol{z}_{t\!-\!1}\!\mid\!\boldsymbol{z}_{t},c\right)d{\boldsymbol{z}_{T}}\!\ldots\!d{\boldsymbol{z}_{1}},$ (1)
where $p\left(\boldsymbol{z}_{t-1}\mid\boldsymbol{z}_{t},c\right)$ denotes the probability of the prior latent representation $\boldsymbol{z}_{t-1}$ based on the current latent representation $\boldsymbol{z}_{t}$ and the user prompt $c$.
Furthermore, by injecting the views condition $\lambda$ into Eq.(1), the latent representations of 2D images with different views are obtained iteratively, as Eq.(2) shows:
$p_{2D}\left(\boldsymbol{z}_{0}\!\mid\!c,\lambda\right)\!=\!\int\!p\left(\boldsymbol{z}_{T}\right)\!\prod_{t=1}^{T}\!p\left(\boldsymbol{z}_{t\!-\!1}\!\mid\!\boldsymbol{z}_{t},c,\lambda\right)d{\boldsymbol{z}_{T}}\!\ldots\!d{\boldsymbol{z}_{1}},$ (2)
The corresponding 3D representation is the joint probability distribution of 2D representations of $N$ iterations. In other words, the multiplication of $n$ latent representation probabilities constructs an unnormalized probability density function for a 3D Gaussian model with parameters $\theta$:
$\tilde{p}_{3D}(\theta)$ $\!=\!\prod_{n=1}^{N}\!\int\!p\left(\boldsymbol{z}_{T}\right)\!\prod_{t=1}^{T}\!p\left(\boldsymbol{z}_{t-1}\!\mid\!\boldsymbol{z}_{t},c,\lambda_{n}\right)d_{\boldsymbol{z}_{T}}\!\ldots\!d_{\boldsymbol{z}_{1}}$ (3)
$=\!\prod_{n=1}^{N}\!p_{2D}\left(\boldsymbol{z}_{0}\!\mid\!c,\lambda_{n}\right),$
where $\lambda_{n}$ is the view of the $n^{\text{th}}$ iteration. It is noticeable that $p_{2D}\left(z_{0}\mid c,\lambda_{n}\right)$ in Eq.(3) illustrates a reverse process given a view $\lambda$ in the diffusion model, and the diffusion model is fixed in training. For emphasis on the procedure from 2D distribution to 3D distribution, Eq.(3) can be simplified as:
$\tilde{p}_{3D}(\theta)=\prod_{n=1}^{N}p\left(Z_{\theta,n}\mid\lambda_{n},c,R_{n}\right),$ (4)
where $R_{n}$ is a 2D rendered image from the 3D content in the ${(n-1)}^{\text{th}}$ iteration, and $Z_{\theta,n}$ denotes the 3D representation at the $n^{\text{th}}$ iteration. In Eq.(4), the user prompt $c$ is expected only to provide content information without views information. However, the current pre-trained T2I models incorporate view prior knowledge when “understanding” user prompt $c$ even if $c$ lacks explicit view information. This prior knowledge arises from the prior preference views of training data. Therefore, we decompose $c$ into a content component $c_{c}$ and a view prior component $c_{v}$, which are expected to be orthogonal:
$\tilde{p}_{3D}(\theta)=\prod_{n=1}^{N}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right),\quad\text{s.t.}\,c_{c}\perp c_{v},$ (5)
[…]
${\rm\log}\tilde{p}_{3D}(\theta)=\sum_{n=1}^{N}{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right),$ (6)
the gradient of $\theta$ denoted as $\nabla\theta{\rm\log}\tilde{p}_{3D}(\theta)$, can be expressed as:
$\nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)=\sum_{n=1}^{N}\nabla_{\theta}{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right)$ (7)
$=\sum_{n=1}^{N}\left(\frac{\partial{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right)}{\partial_{Z_{\theta,n}}}\right)\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},$
as $c_{c}$ and $c_{v}$ are independent, Eq.(7) can be further expanded using Bayes’ theorem. Specifically, the joint distribution $p(Z_{\theta,n},\lambda_{n},c_{c},c_{v},R_{n})$ satisfies the decomposition property, allowing it to be expressed as $p(Z_{\theta,n})p(\lambda_{n},c_{c},c_{v},R_{n}\mid Z_{\theta,n})$. This ensures that the transformation using Bayes’ theorem is valid:
$\nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)\!=\!\sum\limits_{n=1}^{N}\!\left(\!\frac{\partial{\rm\log}p\left(\lambda_{n},c_{c},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\!+\!\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\right)\!\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},\end{aligned}$
where $\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}$ is the unconditional score [33], [49] modelled by T2I models, intuitively reflects the model’s perception of 3D objects without external view guidance. However, as previously mentioned, such perception is prone to prior preference views, which leads to the multi-face Janus problem. To address this issue, Section III-C explores similarity distribution patterns across rendered images and proposes a similarity-based partial order loss $\mathcal{L}_{P}$ to enhance the model’s view awareness in the absence of conditional view information.
Additionally, the second term in Eq.(8) can be further expanded as follows:
$&\frac{\partial}{\partial_{z_{\theta,n}}}{\rm\log}p\left(\lambda_{n},c_{c},c_{v},R_{n}\mid Z_{\theta,n}\right)\\ &\!=\!\underbrace{\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(c_{c},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}_{\text{\normalsize View Bias}}\!+\!\underbrace{\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(\lambda_{n},c_{c},R_{n}\!\mid\!Z_{\theta,n}\right)}_{\text{\normalsize View Control}}\!+\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}M,\end{aligned}$
where $M=\frac{p\left(\lambda_{n},c_{c},c_{v},R_{n}\mid Z_{\theta,n}\right)}{p\left(c_{c},c_{v},R_{n}\mid Z_{\theta,n}\right)p\left(\lambda_{n},c_{c},R_{n}\mid Z_{\theta,n}\right)}$.
$c_{c}$ and $R_{n}$ are constants in the current optimization step, then $M$ can be simplified by using the definition of conditional probability:
$M$ $=\frac{p\left(\lambda_{n},c_{v}\mid Z_{\theta,n}\right)}{p\left(c_{v}\mid Z_{\theta,n}\right)p\left(\lambda_{n}\mid Z_{\theta,n}\right)}$ (10)
$=\frac{p\left(c_{v}\mid Z_{\theta,n}\right)p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)}{p\left(c_{v}\mid Z_{\theta,n}\right)p\left(\lambda_{n}\mid Z_{\theta,n}\right)}$
$=\frac{p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)}{p\left(\lambda_{n}\mid Z_{\theta,n}\right)}.$
Observing Eq.(9) and Eq.(10), when the view control $\lambda_{n}$ and the prior view preference $c_{v}$ conflict under the condition $Z_{\theta,n}$, i.e., $p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)\ll p\left(\lambda_{n}\mid Z_{\theta,n}\right)$, the view bias and the view control in Eq.(9) will have a detrimental impact on the 3D generation jointly. Additionally, in this case, $M$ approaches 0, and the third term $\nabla_{Z_{\theta,n}}{\rm\log}M$ in Eq.(9) introduces a large negative gradient, disrupting the optimization process of the model. Conversely, when $\lambda_{n}$ and $c_{v}$ are consistent, i.e., $p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)\approx p\left(\lambda_{n}\mid Z_{\theta,n}\right)$, the gradient directions of the first two terms in Eq.(9) align, leading to a more stable optimization process. Furthermore, as $M$ approaches 1, the gradient term $\nabla_{Z_{\theta,n}}{\rm\log}M$ approaches 0, ensuring it does not interfere with the optimization process. Therefore, establishing semantic consistency between $c_{v}$ and $\lambda_{n}$ is essential for effectively mitigating the multi-face Janus problem. Section III-B eliminates prior view preferences by disentangling view features and injects precise view control when optimizing the rendering results for specific views, thereby enhancing the semantic consistency between $c_{v}$ and $\lambda_{n}$.
III-B View Disentanglement Module
Based on the mathematical analysis in Section III-A, we aim to eliminate the prior view preference and strengthen the target view specified by the camera parameters. The optimization objective can be expressed as maximizing the consistency between $c_{v}$ and $\lambda_{n}$:
$\min_{\theta}\left|1-M\right|.$ (11)
As $M$ approaches 1, the model effectively understands the target view $\lambda_{n}$ specified by the rendering camera, independent of view biases in the pre-trained model.
Optimizing Eq.(11) is non-trivial, which motivates the proposal of the VDM to eliminate prior view biases and enhance the effectiveness of view control. A two-phase adjustment is applied to achieve the goal: the model’s prior view biases are removed from the original prompt $c$, and the view control is strengthened to ensure the generation process follows the user-specified target view.
The process is illustrated in the upper of Fig.3. First, identify the keyword from the prompt (e.g., the keyword “squirrel” of the prompt “A squirrel eating a burger”), and feed it into the encoder to produce the view-free keyword embedding $\mathbf{V}_{{key}}^{\emptyset}$. Subsequently, new embeddings $\mathbf{V}_{{key}}^{{view}}$ corresponding to various views are obtained by combining the keyword with different view descriptions (e.g., “back view,” “side view,” and “front view”). Unlike $\mathbf{V}_{{key}}^{\phi}$, $\mathbf{V}_{{key}}^{{view}}$ incorporates auxiliary contextual information, which enables the view-specific features $\delta_{{key}}^{{view}}$ isolating content-agnostic information. The process is formulated as follows:
$\delta_{{key}}^{{view}}=\mathbf{V}_{{key}}^{{view}}-{Proj}_{\mathbf{V}_{key}^{\phi}}\left(\mathbf{V}_{{key}}^{{view}}\right),$ (12)
where $Proj_{v}(u)=\frac{u\cdot v}{||v||^{2}}v$ represents the projection of $u$ onto $v$. The view information is deprived from $\mathbf{V}_{{key}}^{{view}}$ by projecting $\mathbf{V}_{{key}}^{{view}}$ to $\mathbf{V}_{{key}}^{\emptyset}$. Furthermore, the view features $\delta_{{key}}^{{view}}$ are obtained by subtracting this projection from the $\mathbf{V}_{{key}}^{{view}}$. The $\delta_{{key}}^{{view}}$ can be applied to user prompts, enabling view control.
In text-to-3D generation tasks, the prior view preferences within the set $\Omega=\{front,side\}$ need to be eliminated first. Subsequently, the view injection process is adaptively implemented using the azimuth $r$ at each optimization step and the intensity coefficient $w$. Specifically, when the azimuth $r$ falls within the range $(-90^{\circ},90^{\circ})$, $\delta_{{key}}^{{side}}$ is injected to suppress prior view preferences from the front view:
[…]
where $w_{1}$ controls the intensity of view injection, $|r|/90$ normalizes the azimuth $r$ symmetrically within the range of -90° to 90°, ensuring that the injection intensity corresponds proportionally to the deviation from the front view. When the azimuth $r$ falls within the range $(-180^{\circ},-90^{\circ})\cup(90^{\circ},180^{\circ})$, both $\delta_{{key}}^{{back}}$ and $\delta_{{key}}^{{side}}$ must be considered simultaneously, because diffusion models may be biased toward generating front views [25]:
[…]
where $w_{2},w_{3}$ control the intensity of view injection. As illustrated in Fig.3, we implement the dynamic adjustment of injection intensity using Eq.(13) and Eq.(14), the model can more accurately “understand” the target view $\lambda_{n}$ specified by the rendering camera. Experimental results in Section IV. demonstrate that this approach effectively mitigates the multi-face Janus problem.
Figure 4: Application of the VDM in 2D Generation. (a) Without the VDM, the model mainly generates front or side views of a car, while the back view is rarely generated. (b) With the VDM, prior view biases are eliminated and targeted view information is injected, allowing the model to successfully generate a car from the back view.
To further evaluate the effectiveness of the VDM, and given that prior view preferences in 3D tasks arise from similar biases in 2D generation, we also apply the VDM to 2D generation tasks. Fig.4 illustrates the impact of VDM on 2D generation tasks. Typically, generated 2D results default to prior preference views. As shown in Fig.4 (a), when the prompt is “A photo of a car, HDR,” the model predominantly generates front or side view images of a car, while images featuring a back view are rarely generated. To generate images from rare views like the “back view,” prior view preferences must be eliminated and target view information injected, as described in Eq.(15):
$\mathbf{V}_{{key}}^{prompt}\!\leftarrow\!\mathbf{V}_{{key}}^{prompt}\underbrace{\!-\!\sum^{\Omega}Proj_{\delta_{{key}}^{{view}}}\!\left(\!\mathbf{V}_{{key}}^{prompt}\!\right)\!}_{\text{\scriptsize Prior Elimination}}\underbrace{+\eta\cdot\delta_{{key}}^{{back}}}_{\text{\scriptsize ControlInjection}},\end{aligned}$
following Eq.(12), the VDM extracts the three view features: $\delta_{car}^{front}$, $\delta_{car}^{side}$, and $\delta_{car}^{back}$. The prior view bias is first eliminated using $\delta_{car}^{front}$ and $\delta_{car}^{side}$, and then the prompt embedding is enhanced by injecting $\delta_{car}^{back}$. As a result, the Stable Diffusion model successfully generates an image of a car from the back view, as shown in Fig.4 (b). This demonstrates the effectiveness of the VDM in overcoming view biases and enabling precise view control in 2D generation tasks.
III-C Partial Order Loss for Cross-View Consistency
The VDM enhances the clarity of view semantics in user prompts, ensuring alignment between user prompts and generated objects in terms of view semantics. However, the mathematical analysis of the multi-face Janus problem in Eq.(8) (Section III-A) reveals that the unconditional term, represented by the gradient $\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial Z_{\theta,n}}$, lacks explicit guidance from view information. This absence of view-specific regulation makes the unconditional term more susceptible to introducing prior view biases into the generated content. During the iterative optimization process, $Z_{\theta,n}$ is projected as a 2D rendered image and input into the Stable Diffusion model, but this 2D projection does not preserve the specific view information in the 3D space. To address this, it is necessary to establish a connection between the unconditional guiding term and view control, ensuring that the unconditional term is regulated by the rendering camera parameters and infusing the model with view awareness.
Figure 5: Partial order loss for cross-view consistency. The random camera views are distributed on the unit circle based on their azimuth angles. A reference camera (e.g., camera 0) is selected, and a red dotted line parallel to the horizontal axis is constructed from this point to mirror the reference to a symmetric position. The azimuthal distances between other camera views and the nearest reference are computed to determine the expected similarity relationships ($sim_{0,3}>sim_{0,2}>sim_{0,1}$). The right-hand side of the figure computes the actual similarity relationships using cosine similarity, and $\mathcal{L}_{P}$ evaluates the alignment between the actual and expected similarity relationships.
Figure 6: Impact of $\mathcal{L}_{P}$ on 3D content over iterations. Using “A squirrel eating a burger” as an example, $\mathcal{L}_{P}$ demonstrates a significant suppression effect on the multi-face Janus problem encountered during early training stages.
To achieve this, we explore the relationship between the multi-view rendered images. A random image is selected from all rendered images as the reference image, and the cosine similarity between the other images and the reference image is computed. As shown in Fig.5, the red camera represents the reference view selected from a series of randomly generated camera views $\Lambda$. It is evident that the closer the azimuth is to the reference azimuth, the higher the similarity score between the corresponding rendered image and the reference image. Furthermore, the scores gradually decrease symmetrically as the azimuth angle deviates from the reference. Therefore, we define a similarity-based partial order loss, where the specific ordering method is illustrated in Fig.5. A Cartesian coordinate system is established, and the random camera views are distributed on the unit circle based on their azimuth angles. One camera view is randomly chosen as the reference view, e.g., camera 0 in Fig.5, and a red dotted line parallel to the horizontal axis is constructed from this point. The reference camera is then mirrored to a symmetric point using this line. The azimuthal distances between other camera views and the nearest reference camera view are computed, determining the expected similarity partial order among the corresponding rendered images: $sim_{0,3}>sim_{0,2}>sim_{0,1}$. The similarity relationship between the 2D projections obtained from Eq.(8) must also adhere to this order. Accordingly, the cosine similarity between the corresponding rendered images is computed as the actual similarity and a similarity-based partial order loss $\mathcal{L}_{P}$ is employed to constrain the actual similarity to align with the expected similarity relationships. $\mathcal{L}_{P}$ is defined as:
$\mathcal{L}_{P}\!=\!\sum_{i=1}^{|\Lambda|-2}\!\max\!\left(0,{sim}\left(R^{i+1},R^{0}\right)\!-\!{sim}\left(R^{i},R^{0}\right)\right),$ (16)
where $R^{i}$ is the 2D rendered image from the 3D content with the $i^{th}$ view in $\Lambda$, $R^{0}$ is the reference image and $\operatorname{sim}\left(R^{i},R^{0}\right)$ denotes the cosine similarity between $R^{i}$ and $R^{0}$. The $\mathcal{L}_{P}$ term can constrain the multi-face Janus problem early in training. As shown in Fig.6, using “A squirrel eating a burger” as an example, without the $\mathcal{L}_{P}$ constraint, the generated result exhibits an irreconcilable multi-face Janus problem. When $\mathcal{L}_{P}$ is applied, the generated 3D content shows significant improvement starting from the 1000th iteration.
Subsequently, we utilize a parameter update module based on DDIM and ISM, which originates from the methodologies outlined in LucidDreamer. Specifically, with a given prompt $c$ and the noisy latents $\boldsymbol{z}_{s}$ and $\boldsymbol{z}_{t}$ generated through DDIM inversion from $\boldsymbol{z}_{0}$ with the time interval $t-s$, the ISM loss is defined as:
$\mathcal{L}_{ISM}(\theta)\!=\!\mathbb{E}_{t,c}\!\left[\omega(t)\!\left\|\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{z}_{t},t,c\right)\!-\!\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{z}_{s},s,\emptyset\right)\right\|_{2}^{2}\right].$ (17)
For detailed information of ISM, please refer to [17]. The final loss function is then formalized as:
${Loss}=\mathcal{L}_{ISM}+\kappa\mathcal{L}_{P},$ (18)
where $\kappa$ is the weight coefficients of $\mathcal{L}_{P}$.
IV Experiments
In this section, we provide a systematic evaluation of ConsDreamer through a dual approach integrating qualitative visual analysis and quantitative benchmarking. First, we conduct a comprehensive ablation study in Section IV-A to evaluate the effectiveness of each individual module. Next, Section IV-B presents comprehensive comparisons against seven SOTA methods to further highlight the advantages of our framework. Finally, we explore the improvement brought by VDM to T2I tasks in Section IV-C, further extending the evaluation. All experiments were conducted using a single NVIDIA RTX 4090 GPU and the Stable Diffusion v2.1
Figure 7: Qualitative results of ablation study. The first row shows the results without the intervention of VDM or $\mathcal{L}_{P}$. The second and third rows display the generated results after applying VDM or $\mathcal{L}_{P}$ respectively. The bottom row presents the performance with both modules jointly optimized. Each generated object is examined from different angles to visually assess its multi-view consistency.
base
Figure 8: Examples of inconsistencies in generated 3D content. The top section highlights the multi-face Janus issue. Additionally, the section below showcases other types of inconsistencies, including component redundancy and component morphing.
model for distillation. Besides, we set the training process with 5,000 iterations and employed the pre-trained Point-E [31] to initialize the 3D Gaussian Splatting.
IV-A Ablation Studies
1) Qualitative Results:
TABLE I: Quantitative Results of Ablation Study on ConsDreamer
[…]
Rank Score Rank Score Rank Score Rank Score Rank Rate Rank Rate
Janus issue 5 -0.770 1 0.340 5 0.086 5 0.050 5 100.0 5 100.0
w/o $\mathcal{L}_{p}$ w/o VDM 4 0.282 5 0.328 3 0.076 4 0.043 4 24.0 4 64.0
w/o VDM 3 0.294 4 0.332 3 0.076 3 0.042 3 15.3 3 52.7
w/o $\mathcal{L}_{p}$ 2 0.465 2 0.333 2 0.068 2 0.039 2 11.4 2 42.0
Ours 1 0.707 2 0.333 1 0.060 1 0.035 1 3.4 1 31.0
  • $\uparrow$
    Higher scores indicate better performance.
  • $\downarrow$
    Lower scores indicate better performance.
By eliminating view semantic bias in T2I mappings and incorporating explicit consistency constraints across multi-view images, our method significantly improves the consistency of generated content. As shown in Fig.7, both VDM and the perceptual loss $\mathcal{L}_{P}$ effectively reduce the occurrence of the multi-face Janus issue. However, certain inconsistency artifacts persist, as demonstrated in the second and third rows of Fig.7, including redundant Batman ears, additional Pikachu limbs, and extraneous nuclear reactors on Iron Man’s back. Notably, when these two modules are combined, the proposed method achieves highly consistent 3D content, surpassing the performance of using either module individually. The results in the bottom row of Fig.7 demonstrate that our approach not only mitigates the multi-face Janus issue but also effectively prevents component redundancy.
Figure 9: Similarity distribution between prompts and rendered results across methods. The similarity distribution between prompts and rendered results for various methods is compared, the results demonstrate that our method generates 3D content with a similarity score distribution that aligns more closely with the ideal distribution.
Figure 10: Qualitative comparisons of 3D content generated by different methods. This figure presents a comparison of our method with seven SOTA text-to-3D generation approaches across various objects, evaluated from five viewpoints (0°, 45°, 90°, 135°, and 180°) for quality and consistency. When generating Super Saiyan characters, DreamFusion, Sherpa3D, LucidDreamer, ScaleDreamer, DreamScene, and GCS-BEG produced redundant limbs. For Iron Man, Sherpa3D, LucidDreamer, ScaleDreamer, and DreamScene incorrectly generated a nuclear reactor on the back. In the case of Batman, Sherpa3D, LucidDreamer, DreamScene, and GCS-BEG generated three ears. The results demonstrate that our proposed ConsDreamer outperforms SOTA methods in terms of both generation quality and multi-view consistency.
TABLE II: Quantitative Comparison with SOTA Methods
[…]
Rank Score Rank Score Rank Score Rank Score Rank Score Rank Rate Rank Rate
DreamFusion [14] 8 3.48 8 4.37 8 3.93 7 -0.289 6 0.307 8 60.0 8 86.7
Debiased-SDS [15] 7 4.19 7 4.93 7 4.56 8 -0.563 8 0.282 4 46.7 2 60.0
Sherpa3D [22] 6 5.42 6 5.82 6 5.62 6 -0.161 7 0.291 7 56.3 7 75.0
LucidDreamer [17] 4 6.76 5 6.31 4 6.54 4 0.285 4 0.328 2 24.0 3 64.0
ScaleDreamer [24] 2 7.82 4 6.33 3 7.08 2 0.371 1 0.333 5 47.3 6 69.2
DreamScene [23] 5 6.25 3 6.90 5 6.58 5 0.201 5 0.321 6 48.3 5 66.7
GCS-BEG [21] 3 7.03 2 7.65 2 7.34 3 0.333 3 0.316 3 33.3 4 64.3
Ours 1 7.85 1 8.05 1 7.95 1 0.707 1 0.333 1 3.4 1 31.0
  • $\uparrow$
    Higher scores indicate better performance.
  • $\downarrow$
    Lower scores indicate better performance.
2) Quantitative Results: Qualitative results visually demonstrate the effectiveness of the proposed framework. We employed conventional evaluation metrics in text-to-3D generation, including ImageReward [50], OpenCLIP-L/14 [51], and A-LPIPS [52], to objectively assess the effectiveness of our framework. ImageReward measures the consistency between generated results and text descriptions, OpenCLIP-L/14 provides enhanced text-image alignment capabilities to capture fine-grained semantic information, and A-LPIPS evaluates the visual quality of generated images, focusing on perceptual similarity. Table I shows that each module of our method outperforms the scenario without intervention in various aspects, and the joint optimization of both modules achieves even better performance. Additionally, we evaluated the percentage of generated results exhibiting Inconsistency, denoted as $f_{inc}$. We broadened the traditional definition of the multi-face Janus problem to encompass a wider range of inconsistencies, such as extra limbs and other implausible artifacts. Fig. 8 provides illustrative examples of Inconsistency. For a comprehensive evaluation, we selected a diverse set of prompts involving objects with distinct front-back differences (e.g., portraits, animals, and vehicles). The quantitative results are summarized in Table I, where the frequency of the multi-face Janus problem is denoted as $f_{mf}$. The results indicate that any of our proposed components significantly improves the consistency of the generated content. Specifically, applying $\mathcal{L}_{p}$ achieves significant improvements, reducing the multi-face problem and enhancing overall consistency by 36.25% and 18.75%, respectively, compared to the scenario without intervention. Similarly, applying VDM demonstrates further enhancements, with improvements of 52.5% and 34.38%, respectively. When both modules are jointly applied, the performance gains are substantially amplified, reaching 85.83% and 51.56%, respectively.
Figure 11: Application of the VDM to T2I Tasks. The results demonstrate the outputs of Stable Diffusion with explicit view descriptions across six prompts. The first column in each set shows the original outputs, highlighting the inherent view bias in T2I models. By applying the VDM, the view features were disentangled, enabling precise control of desired views (e.g., “back view” or “side view”).
However, we chose those generation results with the multi-face Janus problem and observed that the OpenCLIP-L/14 metric yields higher scores in Table I. Fig.9 presents the OpenCLIP similarity distribution between prompts and rendered results for different methods, including samples exhibiting the multi-face Janus problem for comparison. The similarity scores generally decrease progressively from the front view toward both ends, with the lowest scores observed for the back view, which aligns with intuition, as rendered results from the back view typically contain less information, leading to lower similarity between the prompt and the image for that view. However, when the generated object exhibits the multi-face Janus problem, features of prior preference views appear across multiple angles, resulting in abnormally high similarity scores for those views. This observation prompts us to explore more suitable ablation study methods. To more accurately evaluate the consistency of 3D-generated content, analysing the distribution of similarity scores across views is more effective than relying solely on the average similarity score. The results demonstrate that our method produces a similarity score distribution that aligns more closely with the ideal one.
IV-B Comparison with SOTA Methods
1) Qualitative Results: Fig.2 showcases the vivid and diverse text-to-3D results generated by our method. Subsequently, a qualitative comparison is conducted against seven SOTA text-to-3D generation methods. To better assess the quality and view consistency of the generated models, five evenly distributed viewpoints over a 180° azimuth rotation are evaluated. As demonstrated in Fig.10, our method substantially alleviates the multi-face Janus issue prevalent in existing approaches.
2) Quantitative Results: We selected ImageReward, OpenCLIP-L/14, and the frequency of Inconsistency as objective evaluation metrics. As shown in Table II, our method achieves the highest scores across all metrics, particularly in the rate of the multi-face Janus issue, where our method outperforms the average by approximately 90%.
We conducted a user study involving 105 participants to compare with other SOTA methods as subjective evaluation results. Using a 1 to 10 scale, participants assessed the models based on two key criteria: “Generation Quality” and “ 3D Consistency.” To ensure a fair and unbiased comparison, all samples used in the study were randomly selected, with no cherry-picking to favor our approach. As shown in Table II, our method achieves the highest average scores across both evaluation metrics compared to the seven competing methods. These results demonstrate that the 3D content generated by our approach is consistently more appealing and exhibits superior view consistency.
IV-C Application of the VDM to T2I Tasks
1) Qualitative Results: To further validate the effectiveness of the VDM, we extended its application to 2D generation tasks. As shown in Fig.11, we present 18 sets of generation results from the Stable Diffusion model across six prompts, each containing explicit view descriptions (e.g., “back view” or “side view”). The first column in each set shows the original outputs, which reveal the inherent view bias in T2I models. By directly applying the VDM, the view features are effectively disentangled, enabling explicit control over the desired views. The results clearly demonstrate that the VDM mitigates view semantic errors and significantly improves generation accuracy.
2) Quantitative Results: We compared the success rate of generating images with specific views using Stable Diffusion and our method. Given the open-ended nature of the generated content, a strict definition was adopted for counting success cases. Images exhibiting hallucinations (e.g., artifacts or incorrect semantics) were excluded. Additionally, in instances where multiple entities appeared in an image, it was considered a failure if even one entity did not match the specified view. The results, presented in Table III, demonstrate that our method achieves a significantly higher success rate compared to Stable Diffusion, underscoring its effectiveness in aligning multimodal view semantics.
TABLE III: Success Rate of Specific View Generation
Methods Successful Generation Rate(%$\uparrow$
Side view Back view
Stable Diffusion [34] 66.3 28.7
Stable Diffusion+VDM 80.2 77.2
  • $\uparrow$
    Higher scores indicate better performance.
V Conclusion
In this paper, we conducted a mathematical analysis of the multi-face Janus problem in zero-shot text-to-3D generation. To address this challenge, we proposed a novel framework, $ConsDreamer$, which incorporates a View Disentanglement Module to effectively eliminate prior view biases and enhance view control. Additionally, we designed a Partial Order Loss $\mathcal{L}_{P}$ to explicitly enforce multi-view consistency across generated content. We validated the significant improvements of our method over SOTA approaches through comprehensive qualitative and quantitative evaluations. Beyond its application in 3D generation, the View Disentanglement Module also proves effective in 2D generation tasks, correcting view-related semantic biases during the image synthesis process. These contributions collectively establish a robust framework for generating 3D content with greater consistency and realism, paving the way for broader applications in creative industries, virtual reality, and beyond.

### v2

@@
 Abstract
-Recent advances in zero-shot text-to-3D generation have revolutionized 3D content creation by enabling direct synthesis from textual descriptions.
-While state-of-the-art methods leverage 3D Gaussian Splatting with score distillation to enhance multi-view rendering through pre-trained text-to-image (T2I) models, they suffer from inherent view biases in T2I priors.
+Recent advances in zero-shot text-to-3D generation have revolutionised 3D content creation by enabling direct synthesis from textual descriptions.
+While state-of-the-art methods leverage 3D Gaussian Splatting with score distillation to enhance multi-view rendering through pre-trained text-to-image (T2I) models, they suffer from inherent prior view biases in T2I Models.
 These biases lead to inconsistent 3D generation, particularly manifesting as the multi-face Janus problem, where objects exhibit conflicting features across views.
-To address this fundamental challenge, we propose ConsDreamer, a novel framework that mitigates view bias by refining both the conditional and unconditional terms in the score distillation process: (1) a View Disentanglement Module (VDM) that eliminates viewpoint biases in conditional prompts by decoupling irrelevant view components and injecting precise camera parameters; and (2) a similarity-based partial order loss that enforces geometric consistency in the unconditional term by aligning cosine similarities with azimuth relationships.
-Extensive experiments demonstrate that ConsDreamer effectively mitigates the multi-face Janus problem in text-to-3D generation, outperforming existing methods in both visual quality and consistency.
-Index Terms:
-text to 3D generation, multi-face Janus problem, score distillation sampling.
+To address this fundamental challenge, we propose ConsDreamer, a novel method that mitigates view bias by refining both the conditional and unconditional terms in the score distillation process: (1) a View Disentanglement Module (VDM) that eliminates viewpoint biases in conditional prompts by decoupling irrelevant view components and injecting precise view control; and (2) a similarity-based partial order loss that enforces geometric consistency in the unconditional term by aligning cosine similarities with azimuth relationships.
+Extensive experiments demonstrate that ConsDreamer can be seamlessly integrated into various 3D representations and score distillation paradigms, effectively mitigating the multi-face Janus problem.
 publicationid: pubid: Copyright ©20xx IEEE.
@@
 I Introduction
-[…]
-[lines=2]3D GENERATION technology plays a crucial role in various fields such as innovative industrial design, game development, and virtual reality.
-In particular, zero-shot text-to-3D generation [1], [2], [3], [4], [5] aims to generate 3D content without 3D training data, enabling the conversion from concept to reality.
-However, zero-shot text-to-3D generation tasks [6], [7], [8], [9] are constrained by the inherent complexity of the wild world and the scarcity of 3D data, unlike text-to-image (T2I) tasks [10], [11].
+3DGENERATION technology plays a crucial role in various fields such as innovative industrial design, game development, and virtual reality.
+In particular, zero-shot text-to-3D generation [chen2023fantasia3d-lucid-5], [hong2022avatarclip-lucid-14], [lin2023magic3d-lucid-21], [metzer2023latent-lucid-29], [michel2022text2mesh-lucid-31] aims to generate 3D content without 3D training data, enabling the conversion from concept to reality.
+However, zero-shot text-to-3D generation tasks [chen2019text2shape-2405-1], [duan2023dynamic-2405-4], [duan2020sofa-2405-5], [wu2016learning-2405-40] are constrained by the inherent complexity of the wild world and the scarcity of 3D data, unlike text-to-image (T2I) tasks [ruiz2023dreambooth-2405-30], [saharia2022photorealistic-2405-32].
 From this perspective, generating high-quality 3D content from text is still a significant challenge.
-Earlier works [12], [13], [14], [15], [16] based on Neural Radiance Fields (NeRF) introduce a framework that extended the 2D generative capabilities of diffusion models into the 3D domain by leveraging a view-conditioned projection mechanism.
-For instance, DreamFusion [14] introduces the Score Distillation Sampling (SDS) technique, effectively aligning the rendering of 3D content with 2D generative priors.
-However, rendering in NeRF primarily relies on implicit 3D representations, which significantly increases computational cost and makes it challenging to produce high-resolution images [17].
-Recently, 3D Gaussian Splatting [18] has been introduced as an efficient alternative to NeRF’s implicit 3D representation, enabling real-time, high-resolution rendering through explicit point-based modelling.
-Its adoption in SOTA zero-shot text-to-3D generation methods [17], [19], [20], [21] highlights its effectiveness in delivering superior results in quality and speed.
-However, these recent methods still struggle with generation precision, e.g., the multi-face Janus problem [13], [14].
+Figure 1: Examples of the Multi-Face Janus Problem: (a) generated by DreamFusion [poole2022dreamfusion-lucid-34], (b) by SDI [lukoianov2024score], (c) by LucidDreamer [liang2024luciddreamer-lucid], (d) by DreamScene [10.1007/978-3-031-72904-1_13-dreamscene], and (e) by GE3D [li2024text].
+Earlier works [mildenhall2021nerf-lucid-32], [shi2023mvdream-lucid-41], [poole2022dreamfusion-lucid-34], [hong2023debiasing-disbais], [miao2025laser] based on Neural Radiance Fields (NeRF) introduce a framework that extended the 2D generative capabilities of diffusion models into the 3D domain by leveraging a view-conditioned projection mechanism.
+For instance, DreamFusion [poole2022dreamfusion-lucid-34] introduces the Score Distillation Sampling (SDS) technique, effectively aligning the rendering of 3D content with 2D generative priors.
+However, rendering in NeRF primarily relies on implicit 3D representations, which significantly increases computational cost and makes it challenging to produce high-resolution images [liang2024luciddreamer-lucid].
+Recently, 3D Gaussian Splatting [kerbl20233d-2405-13] has been introduced as an efficient alternative to NeRF’s implicit 3D representation, enabling real-time, high-resolution rendering through explicit point-based modelling.
+Its adoption in State-of-the-Art (SOTA) zero-shot text-to-3D generation methods [liang2024luciddreamer-lucid], [sun2024gseditpro-3dgswork1], [di2024hyper-hyper3d], [10.1007/978-3-031-72775-7_16-gcs] highlights its effectiveness in delivering superior results in quality and speed.
+However, these recent methods still struggle with generation precision, e.g., the multi-face Janus problem [shi2023mvdream-lucid-41], [poole2022dreamfusion-lucid-34].
 A single 3D object exhibiting inconsistencies across its faces in different views, as shown in Fig.
@@
 Therefore, ensuring view consistency in 3D generation is crucial for high-quality results, which is the primary motivation for this paper.
-Figure 1: Examples of Multi-Face Janus Problem: (a) generated by DreamFusion [14], (b) by Sherpa3D [22], (c) by LucidDreamer [17], (d) by DreamScene [23], and (e) by ScaleDreamer [24].
-Figure 2: Examples of text-to-3D content creation using our framework.
-We present a text-to-3D generation framework, named $ConsDreamer$, which leverages a View Disentanglement Model and a novel partial order loss to ensure semantic clarity across views (detailed in Section III).
-The generative 3D results demonstrate the superiority of ConsDreamer.
-Please zoom in for details.
+Figure 2: Examples of text-to-3D content creation using our method.
+We introduce ConsDreamerConsDreamer, a method that guarantees multi-view consistency by leveraging a View Disentanglement Model and a novel partial order loss to ensure semantic clarity across views
+(detailed in Section III). The generative 3D results demonstrate the superiority of ConsDreamer. Please zoom in for details.
 Considering the most recent text-to-3D methods, they predominantly rely on T2I models to generate reference images.
@@
 A notable limitation of these datasets is the lack of diverse and well-annotated multi-view images, which results in a bias toward high-frequency single view text-image pairs.
-As a consequence, T2I models tend to prioritise certain viewpoints, particularly frontal or side views [25].
+As a consequence, T2I models tend to prioritise certain viewpoints, particularly frontal or side views [armandpour2023re-perpneg].
 This view bias adversely impacts the consistency of 3D generation.
@@
 Furthermore, the dominance of preferred viewpoints introduces an unbalanced view distribution, where even in the absence of explicit view conditions, the model still favours certain perspectives.
-To tackle the aforementioned problem, we first conducted a mathematical analysis of the Janus problem.
-The analysis reveals that view biases arise during parameter updates, affecting both conditional and unconditional terms used in existing text-to-3D methods.
-In conditional terms, the input prompts guide view control, but conflicts between prior view preferences and target view specifications result in ambiguous view semantics.
-Unconditional terms, independent of prompts, are directly shaped by inherent view biases.
-Therefore, this paper proposes a novel framework, ConsDreamer, to address the challenges from both conditional and unconditional perspectives.
-For the conditional term, we propose the View Disentanglement Module (VDM), which starts extracting view-agnostic keyword embeddings from the prompt by combining the extracted keyword with explicit view descriptions (e.g., “front view,” “side view”).
-These view-specific embeddings are processed to isolate view-related features by subtracting their projection from the view-agnostic embedding space.
-The resulting view-specific features are used to eliminate prior view preferences and inject target view information, ensuring that the generated content adheres to user-specified views.
-For the unconditional term, we introduce a similarity-based partial order loss $\mathcal{L}_{P}$, which leverages the inherent relationship between view similarity and azimuthal angle distances.
+Figure 3: Comparison of PerpNeg and VDM in handling prior view biases.
+(a) Perp-Neg works after UNet denoising in the score space by orthogonalising negative-view scores.
+However, frontal-view features that have already been entangled into the subject-keyword CA map are absorbed into the preserved component $\textit{neg}_{\parallel}$, so prior-view faces survive and still cause Janus artefacts.
+(b) Our VDM instead intervenes before denoising in the prompt-embedding space: it extracts view-specific residuals from view-augmented prompts, subtracts prior-view residuals, and injects only the residual corresponding to the target azimuth, so the conditioning entering the UNet is already view-corrected, producing clean CA maps and substantially reducing Janus.
+To tackle the aforementioned problem, we first conduct a mathematical analysis of the Janus issue.
+This analysis reveals that the prior view biases in T2I models affect both the conditional and unconditional terms used in existing text-to-3D methods.
+The unconditional term, which is independent of the textual prompt, is directly shaped by these inherent view biases.
+In the conditional term, prompt embeddings are fused with the target view description and then fed into the UNet as conditioning signals to generate the specified view.
+However, when the prior view preference of the T2I model disagrees with the desired target view, their interaction leads to ambiguous view semantics and unstable gradients.
+To alleviate this problem, Perp-Neg [armandpour2023re-perpneg] introduces a prompt-handling strategy that orthogonalises view-irrelevant components in the denoising score space, and has been widely adopted in recent text-to-3D pipelines.
+In Perp-Neg, the score corresponding to the target view is treated as the positive condition pos, while the score associated with an irrelevant view is treated as the negative condition neg.
+The method computes an update direction from neg to pos, and decomposes neg into a component parallel to pos, $\textit{neg}_{\parallel}$, and an orthogonal component, $\textit{neg}_{\perp}$.
+The parallel part $\textit{neg}_{\parallel}$ is retained to preserve subject semantics, while $\textit{neg}_{\perp}$ is interpreted as view-specific content and is suppressed.
+However, when the target view conflicts with prior view preference, the cross-attention (CA) map of the subject-keyword (e.g., “squirrel”) still exhibits strong frontal-view activations.
+These erroneous activations are then incorrectly absorbed into $\textit{neg}_{\parallel}$ as subject semantics and thus preserved, which allows the multi-face Janus artefacts to persist.
+The distortions observed in the conditional results closely follow the spatial distribution of these abnormal facial activations, as illustrated in Fig.
+3 (a).
+Therefore, we propose a View Disentanglement Module (VDM) that removes prior view bias and explicitly injects target-view control before the denoising process, thereby avoiding the accumulation of prior-view features inside the UNet.
+VDM computes subject-keyword view-specific residuals from the original prompt and the view-augmented prompt, and the residuals are then used to cancel prior view bias and to inject the desired target-view information according to the camera azimuth.
+In this way, the conditioning representation entering the UNet already aligns with the user-specified view, leading to cleaner CA maps and substantially reducing Janus artefacts in multi-view text-to-3D generation.
+For the unconditional term, we introduce a similarity-based partial order loss $\mathcal{L}_{\text{P}}$, which leverages the inherent relationship between view similarity and azimuthal angle distances.
 To implement this, a Cartesian coordinate framework is established, and an expected similarity partial order is determined based on the proximity of the azimuthal angles of multiple views to a reference view.
-The actual cosine similarity scores of the rendered images are constrained to conform to this partial order.
-By enforcing these constraints, $\mathcal{L}_{P}$ ensures cross-view consistency in 3D content, effectively addressing the multi-face Janus problem.
-We also conducted comprehensive experiments to illustrate the efficacy of our method, confirming that ConsDreamer can better alleviate the multi-face Janus problem and strengthen semantic correlations across different views and enhance generated content quality.
+The actual cosine similarity scores of the rendered-image features are constrained to conform to this partial order.
+By enforcing these constraints, $\mathcal{L}_{\text{P}}$ ensures cross-view consistency in 3D content, effectively addressing the multi-face Janus problem.
+We also conducted comprehensive experiments to illustrate the efficacy of our method, ConsDreamer, which confirms that it can better alleviate the multi-face Janus problem, strengthen inter-view semantic correlations, and enhance the quality of generated content.
 The main contributions of this work are summarized as follows:
@@
   • $\bullet$
-    We conduct extensive experiments on a wide range of prompts to validate the effectiveness of the proposed method.
+    We conduct extensive experiments on a wide range of prompts to verify the effectiveness of the proposed method when integrated into different 3D representations and various score-distillation strategies.
 II Related Works
@@
 Their differentiable nature enables a trainable pathway, optimizing 3D parameters by ensuring that rendered results from random views align with the prompt inputs.
-Previously, various representations are introduced for text-to-3D generation [12], [26], [27], [28].
-Among these, NeRF [12] emerges as the most widely adopted representation in text-to-3D generation tasks.
-NeRF enables novel view synthesis from unobserved viewpoints and leverages guidance information from the T2I model to optimize the tri-plane [29] parameters through a pre-trained MLP network, demonstrating broad applicability across downstream 3D tasks [3], [12], [13], [14], [15].
+Previously, various representations are introduced for text-to-3D generation [mildenhall2021nerf-lucid-32], [barron2021mip-lucid-3], [ge2023ref-lucid-8], [shen2021deep-lucid-40].
+Among these, NeRF [mildenhall2021nerf-lucid-32] emerges as the most widely adopted representation in text-to-3D generation tasks.
+NeRF enables novel view synthesis from unobserved viewpoints and leverages guidance information from the T2I model to optimize the tri-plane [shue20233d-triplane] parameters through a pre-trained MLP network, demonstrating broad applicability across downstream 3D tasks [lin2023magic3d-lucid-21], [mildenhall2021nerf-lucid-32], [shi2023mvdream-lucid-41], [poole2022dreamfusion-lucid-34], [hong2023debiasing-disbais].
 Despite its success, NeRF faces challenges in rendering speed and memory consumption due to its implicit nature.
 A new 3D representation method, named 3D Gaussian Splatting, is thus introduced.
-Gaussian Splatting-based methods [18], [17], [30], [22] address these limitations using explicit 3D representations with anisotropic Gaussians.
+Gaussian Splatting-based methods [kerbl20233d-2405-13], [liang2024luciddreamer-lucid], [tang2023dreamgaussian-lucid-45], [liu2024sherpa3d-sherpa3d] address these limitations using explicit 3D representations with anisotropic Gaussians.
 They ensure high reconstruction quality, real-time rendering, and reduced computational demands, making them ideal for large-scale and dynamic scenes.
-By replacing the NeRF method with the explicit 3D Gaussian Splatting representation, it becomes possible to integrate pre-trained 3D point cloud models like Point-E [31] and Shape-E [32] into the 3D generation framework, providing a prior 3D knowledge foundation.
-II-B Diffusion models
-Diffusion models serve as a core component for guiding text-to-3D generation.
-The Denoising Diffusion Probabilistic Model (DDPM) [33], based on Markov chains, operates by gradually “diffusing” data from its original state into random noise in the forward procedure and then reversing the process to reconstruct the original data.
-Although DDPM achieves “diffusion-reconstruction,” its efficiency is hindered by the step-by-step sampling strategy.
-Subsequently, Denoising Diffusion Implicit Models (DDIM) [11] introduce a deterministic sampling approach, enabling high-quality generation with significantly fewer sampling steps.
-Additionally, DDPM faces high time consumption as the data retains its original size at every step.
-Latent Diffusion Models (LDM) [34] address this limitation by transferring the diffusion process from the original pixel space to a low-dimensional latent space, significantly improving efficiency, where the attention [7], [35] mechanism could help to control the learning process.
-Most recently, Stable Diffusion [34] introduces a conditional generation approach based on the LDM, enabling high-quality image generation from textual descriptions.
-Subsequent research [17], [36], [37] extends Stable Diffusion with the deterministic sampling mechanism in DDIM.
-II-C Text-to-3D Generation
-Early efforts, such as DreamField [38], leverage the cross-modal information of CLIP [39] to translate text into 3D.
+By replacing the NeRF method with the explicit 3D Gaussian Splatting representation, it becomes possible to integrate pre-trained 3D point cloud models like Point-E [nichol2022point-lucid-33] and Shape-E [jun2023shap-sherpa3d-28] into the 3D generation framework, providing a prior 3D knowledge foundation.
+II-B Text-to-3D Generation
+Early efforts, such as DreamField [jain2022zero-2405-10], leverage the cross-modal information of CLIP [ramesh2021zero-2405-26] to translate text into 3D.
 However, DreamField faces limitations due to the lack of high-precision alignment in CLIP’s cross-modal information, resulting in suboptimal 3D generation performance.
-With the advent of diffusion models, a new paradigm [14], [30] emerges for extracting 3D assets from pre-trained T2I models, leveraging 2D diffusion model outputs to guide differentiable 3D representations.
+With the advent of diffusion models, a new paradigm [poole2022dreamfusion-lucid-34], [tang2023dreamgaussian-lucid-45] emerges for extracting 3D assets from pre-trained T2I models, leveraging 2D diffusion model outputs to guide differentiable 3D representations.
 For example, DreamFusion introduces Score Distillation Sampling (SDS), which aligns 3D rendering information with diffusion priors and optimizes model parameters through backpropagation.
 However, SDS-based approaches often introduce smoothing effects.
-Recent research [40], [41], [42], [43] aims to address this.
-Prolific-Dreamer [41] proposes Variational Score Distillation (VSD), significantly improving generation quality.
-Consistent 3D [9] analyzes SDS from ordinary differential equations (ODEs) and introduces Consistency Distillation Sampling (CSD).
-ScaleDreamer [24] introduces Asynchronous Score Distillation (ASD), which shifts diffusion timesteps to minimize noise prediction errors without fine-tuning the diffusion model.
-LucidDreamer [17] addresses over-smoothing issues via Interval Score Matching (ISM), integrating DDIM inversion to enable deterministic sampling.
-Subsequent advancements, DreamerXL [37], ExactDreamer [36], and Guided Consistency Sampling (GCS) [21] further refine this paradigm by systematically reducing accumulated errors in DDIM’s reverse process.
-The methods above are predominantly optimization-based, while recent advances also explore feed-forward methodologies [44], [45], [46] to facilitate the generation of large-scale 3D assets efficiently [47] or to integrate auxiliary information such as text-to-3D knowledge graphs [48].
+Recent research [katzir2023noise-lucid-18], [wang2024prolificdreamer-lucid-47], [yu2023text-lucid-49], [zhu2023hifa-lucid-52] aims to address this.
+Prolific-Dreamer [wang2024prolificdreamer-lucid-47] proposes Variational Score Distillation (VSD), significantly improving generation quality.
+Consistent 3D [wu2016learning-2405-40] analyzes SDS from ordinary differential equations (ODEs) and introduces Consistency Distillation Sampling (CSD).
+ScaleDreamer [10.1007/978-3-031-72667-5_1-scaledreamer] introduces Asynchronous Score Distillation (ASD), which shifts diffusion timesteps to minimize noise prediction errors without fine-tuning the diffusion model.
+LucidDreamer [liang2024luciddreamer-lucid] addresses over-smoothing issues via Interval Score Matching (ISM), integrating Denoising Diffusion Implicit Models (DDIM) [saharia2022photorealistic-2405-32] inversion to enable deterministic sampling.
+Subsequent advancements, DreamerXL [miao2024dreamer-dreamxl], ExactDreamer [zhang2024exactdreamer-exactdreamer], and Guided Consistency Sampling (GCS) [10.1007/978-3-031-72775-7_16-gcs] further refine this paradigm by systematically reducing accumulated errors in DDIM’s reverse process.
+The methods above are predominantly optimization-based, while recent advances also explore feed-forward methodologies [xu2023dmv3d], [liu2024pi3d], [cao2023large-difftf] to facilitate the generation of large-scale 3D assets efficiently [cao2025difftf++] or to integrate auxiliary information such as text-to-3D knowledge graphs [nie2024t2td-t2td].
 These methods, however, demand substantial GPU resources for additional training and require extensive 3D datasets.
+II-C Mitigating the Multi-Face Janus Issue
+Several studies have specifically tackled the multi-face Janus problem.
+Among these, Perp-Neg [armandpour2023re-perpneg] stands out as the most widely adopted, which leverages the geometric properties of the score space to enhance the utilisation of negative prompts.
+However, this method operates solely on the UNet output, thereby neglecting the highly aggregated prior view information throughout the UNet processing pipeline.
+Consequently, the multi-face Janus problem may still recur when lifting 2D representations to 3D.
+MVDream [shi2023mvdream-lucid-41], Zero123 [liu2023zero1to3], and SV3D [voleti2024sv3d] all rely on pre-trained models and finetuning, which introduces high computational costs and dependency on large datasets.
+While MVDream ensures view consistency through fine-tuning, and Zero123 and SV3D generate multi-view 3D representations, these methods often struggle with maintaining consistent 3D geometry and texture across views.
+Moreover, they require extensive training, making them less efficient for quick deployment.
+In this paper, we introduce a novel method that eliminates prior view biases by disentangling view features and ensuring multi-view consistency through a carefully designed partial order loss.
+This approach significantly enhances view semantic clarity without the need for additional training or 3D datasets, offering a versatile, rapid-deployment solution that can be easily applied to a wide range of 3D generation frameworks.
 Figure 4: An overview of ConsDreamer.
-Our framework is built upon the main flow of 3D content distilled from the T2I model.
+Our method is built upon the main flow of 3D content distilled from the T2I model.
 ConsDreamer introduces two key innovations: (a) VDM disentangles the keyword in the prompt to obtain the canonical view features $\delta_{\text{key}}^{\text{view}}$, which are then used for precise view control (detailed in Section III-B).
-(b) A novel partial order loss $\mathcal{L}_{P}$ is introduced among multi-view rendered images to endow the model with view-aware capabilities (detailed in Section III-C).
-Together, the VDM and $\mathcal{L}_{P}$ enhance the clarity of view semantics and significantly mitigate the multi-face Janus problem.
-Several works specifically address the multi-face Janus problem.
-Perp-Neg [25] leverages the geometric properties of score space to enhance the utilization of negative prompts.
-MVDream [13] fine-tunes pre-trained T2I models using canonical view image combinations, ensuring view consistency but requiring additional training and datasets.
-Debiased-SDS [15] mitigates the multi-face Janus problem in two ways: first, by clipping distillation scores to suppress erroneous artifacts, and second, by identifying and omitting words in the prompt that conflict with the target view.
-However, Debiased-SDS lacks fine-grained view control, and omitting conflicting words can compromise the integrity of user prompts.
-Sherpa3D [22] uses structural alignment and semantic guidance, incorporating normal map gradients and CLIP semantic embeddings to preserve the geometric consistency of the initial 3D prior during 2D lifting optimization.
-However, this method heavily relies on coarse 3D prior information.
-Overall, existing approaches lack precise understanding and control of view-related information while requiring additional computational resources.
-In this paper, we introduce a novel framework that eliminates prior view biases by disentangling view features and ensuring multi-view consistency through a carefully designed partial order loss, significantly enhancing view semantic clarity without requiring additional training.
+(b) A novel partial order loss $\mathcal{L}_{\text{P}}$ is introduced among multi-view rendered images to endow the model with view-aware capabilities (detailed in Section III-C).
+Together, the VDM and $\mathcal{L}_{\text{P}}$ enhance the clarity of view semantics and significantly mitigate the multi-face Janus problem.
 III Methodology
-To tackle the multi-face Janus problem, we propose ConsDreamer, a unified framework to enhance view semantic clarity.
+TABLE I: Summary of main symbols and notations used in the analysis.
+Symbol Description
+$c$ User-provided text prompt.
+$c_{sbj}$ Subject-related component of $c$.
+$c_{v}$ Prior view bias component of $c$.
+$\lambda$ View condition in diffusion model.
+$\lambda_{n}$ View condition of $n$-th optimization step.
+$r$ Azimuth angle of camera for $\lambda_{n}$.
+$\Omega$ Set of views with strong priors, $\Omega=\{\text{front},\text{side}\}$.
+$R_{n}$ 2D image rendered from 3D representation at step n1n\!-\!1.
+$\mathcal{L}_{\text{P}}$ Similarity-based partial order loss.
+$\boldsymbol{z}_{t}$ 2D latent at diffusion step $t$, $t\in\{0,\ldots,T\}$.
+$\boldsymbol{z}_{0}$ Clean 2D latent (reverse diffusion output).
+$\boldsymbol{z}_{T}$ Final noise latent sampled from $p(\boldsymbol{z}_{T})$.
+$T$ Total diffusion time steps.
+$p(\cdot)$ Probability density function.
+$p_{2D}(\boldsymbol{z}_{0}\mid c)$ 2D latent distribution (conditioned on $c$).
+$\tilde{p}_{3D}(\theta)$ Unnormalized PDF of 3D representation.
+NN Number of optimisation iterations.
+$\theta$ 3D parameters.
+$Z_{\theta,n}$ 3D representation at iteration $n$ (parameterized by $\theta$).
+$\nabla_{\theta}$ Gradient operator w.r.t. $\theta$.
+$M$ Compatibility term ($c_{v}$ vs. $\lambda_{n}$, Eq. (11)).
+$\langle\cdot,\cdot\rangle$ Inner product of two vectors.
+$\|\cdot\|$ Euclidean norm of a vector.
+$|\cdot|$ Absolute value of a scalar.
+$\perp$ Orthogonality (e.g., $c_{sbj}\perp c_{v}$).
+$\mathbf{V}^{\emptyset}_{\text{key}}$ Keyword embedding (no view description).
+$\mathbf{V}^{\text{view}}_{\text{key}}$ Keyword embedding (with specific view description).
+$\delta^{\text{view}}_{\text{key}}$ View-specific residual feature.
+$Proj_{v}(u)$ Projection of $u$ onto $v$: $Proj_{v}(u)=\frac{u\cdot v}{\|v\|^{2}}v$.
+To tackle the multi-face Janus problem, we propose ConsDreamer, a unified method to enhance view semantic clarity.
 Fig.4 provides an overview of our approach.
@@
 Based on this analysis, Section III-B introduces VDM to eliminate prior view preferences and integrate target view control in the conditional term.
-Meanwhile, for the unconditional term, Section III-C explores similarity distribution patterns across views in rendered images and proposes a similarity-based partial order loss $\mathcal{L}_{P}$ to enhance spatial consistency in 3D content.
+Meanwhile, for the unconditional term, Section III-C explores similarity distribution patterns across views in rendered images and proposes a similarity-based partial order loss $\mathcal{L}_{\text{P}}$ to enhance spatial consistency in 3D content.
+For clarity, the main symbols and notations used in this section are summarised in Table I.
 III-A Analysis of the multi-face Janus problem
@@
 where $R_{n}$ is a 2D rendered image from the 3D content in the ${(n-1)}^{\text{th}}$ iteration, and $Z_{\theta,n}$ denotes the 3D representation at the $n^{\text{th}}$ iteration.
-In Eq.(4), the user prompt $c$ is expected only to provide content information without views information.
+In Eq.(4), the user prompt $c$ is expected only to provide subject information without views information.
 However, the current pre-trained T2I models incorporate view prior knowledge when “understanding” user prompt $c$ even if $c$ lacks explicit view information.
 This prior knowledge arises from the prior preference views of training data.
-Therefore, we decompose $c$ into a content component $c_{c}$ and a view prior component $c_{v}$, which are expected to be orthogonal:
-$\tilde{p}_{3D}(\theta)=\prod_{n=1}^{N}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right),\quad\text{s.t.}\,c_{c}\perp c_{v},$ (5)
+Therefore, we decompose $c$ into a subject component $c_{sbj}$ and a view prior component $c_{v}$, which are expected to be orthogonal:
+$\tilde{p}_{3D}(\theta)=\prod_{n=1}^{N}p\left(Z_{\theta,n}\mid\lambda_{n},c_{sbj},c_{v},R_{n}\right),\quad\text{s.t.}\,c_{sbj}\perp c_{v},$ (5)
 […]
-${\rm\log}\tilde{p}_{3D}(\theta)=\sum_{n=1}^{N}{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right),$ (6)
-the gradient of $\theta$ denoted as $\nabla\theta{\rm\log}\tilde{p}_{3D}(\theta)$, can be expressed as:
-$\nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)=\sum_{n=1}^{N}\nabla_{\theta}{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right)$ (7)
-$=\sum_{n=1}^{N}\left(\frac{\partial{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right)}{\partial_{Z_{\theta,n}}}\right)\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},$
-as $c_{c}$ and $c_{v}$ are independent, Eq.(7) can be further expanded using Bayes’ theorem.
-Specifically, the joint distribution $p(Z_{\theta,n},\lambda_{n},c_{c},c_{v},R_{n})$ satisfies the decomposition property, allowing it to be expressed as $p(Z_{\theta,n})p(\lambda_{n},c_{c},c_{v},R_{n}\mid Z_{\theta,n})$.
+${\rm\log}\tilde{p}_{3D}(\theta)=\sum_{n=1}^{N}{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{sbj},c_{v},R_{n}\right),$ (6)
+the gradient of $\theta$ denoted as $\nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)$, can be expressed as:
+$\nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)=\sum_{n=1}^{N}\nabla_{\theta}{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{sbj},c_{v},R_{n}\right)$ (7)
+$=\sum_{n=1}^{N}\left(\frac{\partial{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{sbj},c_{v},R_{n}\right)}{\partial_{Z_{\theta,n}}}\right)\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},$
+as $c_{sbj}$ and $c_{v}$ are independent, Eq.(7) can be further expanded using Bayes’ theorem.
+Specifically, the joint distribution $p(Z_{\theta,n},\lambda_{n},c_{sbj},c_{v},R_{n})$ satisfies the decomposition property, allowing it to be expressed as $p(Z_{\theta,n})p(\lambda_{n},c_{sbj},c_{v},R_{n}\mid Z_{\theta,n})$.
 This ensures that the transformation using Bayes’ theorem is valid:
-$\nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)\!=\!\sum\limits_{n=1}^{N}\!\left(\!\frac{\partial{\rm\log}p\left(\lambda_{n},c_{c},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\!+\!\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\right)\!\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},\end{aligned}$
-where $\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}$ is the unconditional score [33], [49] modelled by T2I models, intuitively reflects the model’s perception of 3D objects without external view guidance.
+$\begin{aligned} \nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)&=\sum\limits_{n=1}^{N}\!\left(\!\frac{\partial{\rm\log}p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\right.\\ &\quad\left.+\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\!-\!\frac{\partial{\rm\log}p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\right)}{\partial_{Z_{\theta,n}}}\right)\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},\end{aligned}$
+the term $\frac{\partial{\rm\log}p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\right)}{\partial_{Z_{\theta,n}}}$ is a constant term independent of $Z_{\theta,n}$, and its derivative is zero, thus can be removed:
+$\begin{aligned} \nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)\!=\!\sum\limits_{n=1}^{N}\!\left(\!\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}+\!\frac{\partial{\rm\log}p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\!\right)\!\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},\end{aligned}$
+where $\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}$ is the unconditional score [ho2020denoising-2405-7], [song2020score-2405-35] modelled by T2I models, intuitively reflects the model’s perception of 3D objects without external view guidance.
 However, as previously mentioned, such perception is prone to prior preference views, which leads to the multi-face Janus problem.
-To address this issue, Section III-C explores similarity distribution patterns across rendered images and proposes a similarity-based partial order loss $\mathcal{L}_{P}$ to enhance the model’s view awareness in the absence of conditional view information.
+To address this issue, Section III-C explores similarity distribution patterns across rendered images and proposes a similarity-based partial order loss $\mathcal{L}_{\text{P}}$ to enhance the model’s view awareness in the absence of conditional view information.
 Additionally, the second term in Eq.(9) can be further expanded as follows:
-$&\frac{\partial}{\partial_{z_{\theta,n}}}{\rm\log}p\left(\lambda_{n},c_{c},c_{v},R_{n}\mid Z_{\theta,n}\right)\\ &\!=\!\underbrace{\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(c_{c},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}_{\text{\normalsize View Bias}}\!+\!\underbrace{\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(\lambda_{n},c_{c},R_{n}\!\mid\!Z_{\theta,n}\right)}_{\text{\normalsize View Control}}\!+\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}M,\end{aligned}$
-where $M=\frac{p\left(\lambda_{n},c_{c},c_{v},R_{n}\mid Z_{\theta,n}\right)}{p\left(c_{c},c_{v},R_{n}\mid Z_{\theta,n}\right)p\left(\lambda_{n},c_{c},R_{n}\mid Z_{\theta,n}\right)}$.
-$c_{c}$ and $R_{n}$ are constants in the current optimization step, then $M$ can be simplified by using the definition of conditional probability:
+$\begin{aligned} &\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\mid Z_{\theta,n}\right)\\ &\!=\!\underbrace{\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(c_{sbj},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}_{\text{\normalsize Prior View Bias}}\!+\!\underbrace{\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(\lambda_{n},c_{sbj},R_{n}\!\mid\!Z_{\theta,n}\right)}_{\text{\normalsize Target View Control}}\!+\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}M,\end{aligned}$
+where $\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(c_{sbj},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)$ is independent of the view condition $\lambda_{n}$ and denotes the prior view bias term, while $\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(\lambda_{n},c_{sbj},R_{n}\!\mid\!Z_{\theta,n}\right)$ encodes view information exclusively governed by the view condition $\lambda_{n}$, corresponding to the target view control term, and $M=~\frac{p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\mid Z_{\theta,n}\right)}{p\left(c_{sbj},c_{v},R_{n}\mid Z_{\theta,n}\right)p\left(\lambda_{n},c_{sbj},R_{n}\mid Z_{\theta,n}\right)}$.
+$c_{sbj}$ and $R_{n}$ are constants in the currentoptimisationn step, then $M$ can be simplified by using the definition of conditional probability:
 $M$ $=\frac{p\left(\lambda_{n},c_{v}\mid Z_{\theta,n}\right)}{p\left(c_{v}\mid Z_{\theta,n}\right)p\left(\lambda_{n}\mid Z_{\theta,n}\right)}$ (11)
@@
 $=\frac{p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)}{p\left(\lambda_{n}\mid Z_{\theta,n}\right)}.$
-Observing Eq.(9) and Eq.(10), when the view control $\lambda_{n}$ and the prior view preference $c_{v}$ conflict under the condition $Z_{\theta,n}$, i.e., $p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)\ll p\left(\lambda_{n}\mid Z_{\theta,n}\right)$, the view bias and the view control in Eq.(9) will have a detrimental impact on the 3D generation jointly.
-Additionally, in this case, $M$ approaches 0, and the third term $\nabla_{Z_{\theta,n}}{\rm\log}M$ in Eq.(9) introduces a large negative gradient, disrupting the optimization process of the model.
+Observing Eq.(10) and Eq.(11), when the view control $\lambda_{n}$ and the prior view preference $c_{v}$ conflict under the condition $Z_{\theta,n}$, i.e., $p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)\ll p\left(\lambda_{n}\mid Z_{\theta,n}\right)$, the prior view bias and the target view control in Eq.(10) will have a detrimental impact on the 3D generation jointly, which echoes the abnormal activations observed in the subject-keyword CA map in Fig.
+3(a).
+Additionally, in this case, $M$ approaches 0, and the third term $\nabla_{Z_{\theta,n}}{\rm\log}M$ in Eq.(10) introduces a large negative gradient, disrupting the optimisation process of the model.
 Conversely, when $\lambda_{n}$ and $c_{v}$ are consistent, i.e., $p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)\approx p\left(\lambda_{n}\mid Z_{\theta,n}\right)$, the gradient directions of the first two terms in Eq.(10) align, leading to a more stable optimization process.
-Furthermore, as $M$ approaches 1, the gradient term $\nabla_{Z_{\theta,n}}{\rm\log}M$ approaches 0, ensuring it does not interfere with the optimization process.
+Furthermore, as $M$ approaches 1, the gradient term $\nabla_{Z_{\theta,n}}{\rm\log}M$ approaches 0, ensuring it does not interfere with the optimisation process.
 Therefore, establishing semantic consistency between $c_{v}$ and $\lambda_{n}$ is essential for effectively mitigating the multi-face Janus problem.
-Section III-B eliminates prior view preferences by disentangling view features and injects precise view control when optimizing the rendering results for specific views, thereby enhancing the semantic consistency between $c_{v}$ and $\lambda_{n}$.
+Section III-B eliminates prior view preferences by disentangling view features and injects precise view control when optimising the rendering results for specific views, thereby enhancing the semantic consistency between $c_{v}$ and $\lambda_{n}$.
 III-B View Disentanglement Module
 Based on the mathematical analysis in Section III-A, we aim to eliminate the prior view preference and strengthen the target view specified by the camera parameters.
-The optimization objective can be expressed as maximizing the consistency between $c_{v}$ and $\lambda_{n}$:
+The optimisation objective can be expressed as maximising the consistency between $c_{v}$ and $\lambda_{n}$:
 $\min_{\theta}\left|1-M\right|.$ (12)
@@
 The process is illustrated in the upper of Fig.4.
-First, identify the keyword from the prompt (e.g., the keyword “squirrel” of the prompt “A squirrel eating a burger”), and feed it into the encoder to produce the view-free keyword embedding $\mathbf{V}_{{key}}^{\emptyset}$.
-Subsequently, new embeddings $\mathbf{V}_{{key}}^{{view}}$ corresponding to various views are obtained by combining the keyword with different view descriptions (e.g., “back view,” “side view,” and “front view”).
-Unlike $\mathbf{V}_{{key}}^{\phi}$, $\mathbf{V}_{{key}}^{{view}}$ incorporates auxiliary contextual information, which enables the view-specific features $\delta_{{key}}^{{view}}$ isolating content-agnostic information.
+First, identify the keyword from the prompt (e.g., the keyword “squirrel” of the prompt “A squirrel eating a burger”), and extract the keyword embedding $\mathbf{V}_{{key}}^{\emptyset}$ from the prompt encoding results.
+Subsequently, new embeddings $\mathbf{V}_{{key}}^{{view}}$ corresponding to various views are obtained by combining the prompt with different view descriptions (e.g., “back view,” “side view,” and “front view”).
+Unlike $\mathbf{V}_{{key}}^{\emptyset}$, $\mathbf{V}_{{key}}^{{view}}$ incorporates auxiliary contextual information, which enables the view-specific features $\delta_{{key}}^{{view}}$ isolating content-agnostic information.
 The process is formulated as follows:
-$\delta_{{key}}^{{view}}=\mathbf{V}_{{key}}^{{view}}-{Proj}_{\mathbf{V}_{key}^{\phi}}\left(\mathbf{V}_{{key}}^{{view}}\right),$ (12)
-where $Proj_{v}(u)=\frac{u\cdot v}{||v||^{2}}v$ represents the projection of $u$ onto $v$.
-The view information is deprived from $\mathbf{V}_{{key}}^{{view}}$ by projecting $\mathbf{V}_{{key}}^{{view}}$ to $\mathbf{V}_{{key}}^{\emptyset}$.
-Furthermore, the view features $\delta_{{key}}^{{view}}$ are obtained by subtracting this projection from the $\mathbf{V}_{{key}}^{{view}}$.
-The $\delta_{{key}}^{{view}}$ can be applied to user prompts, enabling view control.
+$\delta_{{key}}^{{view}}=\mathbf{V}_{{key}}^{{view}}-{Proj}_{\mathbf{V}_{key}^{\emptyset}}\left(\mathbf{V}_{{key}}^{{view}}\right),$ (13)
+[…]
 In text-to-3D generation tasks, the prior view preferences within the set $\Omega=\{front,side\}$ need to be eliminated first.
@@
 Specifically, when the azimuth $r$ falls within the range $(-90^{\circ},90^{\circ})$, $\delta_{{key}}^{{side}}$ is injected to suppress prior view preferences from the front view:
-[…]
-where $w_{1}$ controls the intensity of view injection, $|r|/90$ normalizes the azimuth $r$ symmetrically within the range of -90° to 90°, ensuring that the injection intensity corresponds proportionally to the deviation from the front view.
-When the azimuth $r$ falls within the range $(-180^{\circ},-90^{\circ})\cup(90^{\circ},180^{\circ})$, both $\delta_{{key}}^{{back}}$ and $\delta_{{key}}^{{side}}$ must be considered simultaneously, because diffusion models may be biased toward generating front views [25]:
-[…]
-[…]
-where $w_{2},w_{3}$ control the intensity of view injection.
-As illustrated in Fig.3, we implement the dynamic adjustment of injection intensity using Eq.(13) and Eq.(14), the model can more accurately “understand” the target view $\lambda_{n}$ specified by the rendering camera.
-Experimental results in Section IV.
-demonstrate that this approach effectively mitigates the multi-face Janus problem.
+$\mathbf{V}^{\emptyset}_{\mathit{key}}\!\leftarrow\!\mathbf{V}^{\emptyset}_{\mathit{key}}\!-\!\sum^{\Omega}\!Proj_{\delta_{{key}}^{{view}}}(\mathbf{V}^{\emptyset}_{\mathit{key}})\!+\!w_{1}\!\cdot\!\frac{|r|}{90}\!\cdot\!\delta^{\mathit{side}}_{\mathit{key}},$ (14)
 Figure 5: Application of the VDM in 2D Generation.
@@
 (b) With the VDM, prior view biases are eliminated and targeted view information is injected, allowing the model to successfully generate a car from the back view.
+where $w_{1}$ controls the intensity of view injection, $|r|/90$ normalizes the azimuth $r$ symmetrically within the range of -90° to 90°, ensuring that the injection intensity corresponds proportionally to the deviation from the front view.
+When the azimuth $r$ falls within the range $(-180^{\circ},-90^{\circ})\cup(90^{\circ},180^{\circ})$, both $\delta_{{key}}^{{back}}$ and $\delta_{{key}}^{{side}}$ must be considered simultaneously, because diffusion models may be biased toward generating front views [armandpour2023re-perpneg]:
+$\mathbf{V}^{\emptyset}_{\mathit{key}}\!\leftarrow$ $\mathbf{V}^{\emptyset}_{\mathit{key}}\!-\!\sum^{\Omega}Proj_{\delta_{{key}}^{{view}}}(\mathbf{V}^{\emptyset}_{\mathit{key}})$ (15)
+$\!+\!w_{2}\!\cdot\!\frac{|r|\!-\!90}{90}\!\cdot\!\delta^{\mathit{back}}_{\mathit{key}}\!+\!w_{3}\cdot\frac{180\!-\!|r|}{90}\!\cdot\!\delta^{\mathit{side}}_{\mathit{key}},$
+where $w_{2},w_{3}$ control the intensity of view injection.
+As illustrated in Fig.4, we implement the dynamic adjustment of injection intensity using Eq.(14) and Eq.(15), the model can more accurately “understand” the target view $\lambda_{n}$ specified by the rendering camera.
+We fix $w_{1}=1.0$, $w_{2}=1.5$ and $w_{3}=1.0$ in all experiments, and results in Section IV demonstrate that this approach effectively mitigates the multi-face Janus problem.
+Figure 6: Partial order loss for cross-view consistency.
+The random camera views are distributed on the unit circle based on their azimuth angles.
+A reference camera (e.g., camera 0) is selected, and a red dotted line parallel to the horizontal axis is constructed from this point to mirror the reference to a symmetric position.
+The azimuthal distances between other camera views and the nearest reference are computed to determine the expected similarity relationships ($\text{sim}_{0,3}>\text{sim}_{0,2}>\text{sim}_{0,1}$).
+The right-hand side of the figure computes the actual similarity relationships using cosine similarity, and $\mathcal{L}_{\text{P}}$ evaluates the alignment between the actual and expected similarity relationships.
 To further evaluate the effectiveness of the VDM, and given that prior view preferences in 3D tasks arise from similar biases in 2D generation, we also apply the VDM to 2D generation tasks.
@@
 To generate images from rare views like the “back view,” prior view preferences must be eliminated and target view information injected, as described in Eq.(16):
-$\mathbf{V}_{{key}}^{prompt}\!\leftarrow\!\mathbf{V}_{{key}}^{prompt}\underbrace{\!-\!\sum^{\Omega}Proj_{\delta_{{key}}^{{view}}}\!\left(\!\mathbf{V}_{{key}}^{prompt}\!\right)\!}_{\text{\scriptsize Prior Elimination}}\underbrace{+\eta\cdot\delta_{{key}}^{{back}}}_{\text{\scriptsize ControlInjection}},\end{aligned}$
+$\begin{aligned} \mathbf{V}_{{key}}^{prompt}\!\leftarrow\!\mathbf{V}_{{key}}^{prompt}\underbrace{\!-\!\sum^{\Omega}Proj_{\delta_{{key}}^{{view}}}\!\left(\!\mathbf{V}_{{key}}^{prompt}\!\right)\!}_{\text{\scriptsize Prior Elimination}}\underbrace{+\eta\cdot\delta_{{key}}^{{back}}}_{\text{\scriptsize Control Injection}},\end{aligned}$
 following Eq.(13), the VDM extracts the three view features: $\delta_{car}^{front}$, $\delta_{car}^{side}$, and $\delta_{car}^{back}$.
@@
 This demonstrates the effectiveness of the VDM in overcoming view biases and enabling precise view control in 2D generation tasks.
+Figure 7: Impact of $\mathcal{L}_{\text{P}}$ on 3D content over iterations.
+Using “A squirrel eating a burger” as an example, $\mathcal{L}_{\text{P}}$ demonstrates a significant suppression effect on the multi-face Janus problem encountered during early training stages.
 III-C Partial Order Loss for Cross-View Consistency
@@
 To address this, it is necessary to establish a connection between the unconditional guiding term and view control, ensuring that the unconditional term is regulated by the rendering camera parameters and infusing the model with view awareness.
-Figure 5: Partial order loss for cross-view consistency.
-The random camera views are distributed on the unit circle based on their azimuth angles.
-A reference camera (e.g., camera 0) is selected, and a red dotted line parallel to the horizontal axis is constructed from this point to mirror the reference to a symmetric position.
-The azimuthal distances between other camera views and the nearest reference are computed to determine the expected similarity relationships ($sim_{0,3}>sim_{0,2}>sim_{0,1}$).
-The right-hand side of the figure computes the actual similarity relationships using cosine similarity, and $\mathcal{L}_{P}$ evaluates the alignment between the actual and expected similarity relationships.
-Figure 6: Impact of $\mathcal{L}_{P}$ on 3D content over iterations.
-Using “A squirrel eating a burger” as an example, $\mathcal{L}_{P}$ demonstrates a significant suppression effect on the multi-face Janus problem encountered during early training stages.
+Figure 8: Qualitative ablation study results (prompt: “A photo of a Pikachu, yellow, electric”).
+Incrementally integrating our proposed method’s components into the baseline enhances generation consistency.
+To enhance the persuasiveness of our analysis, we include Perp-Neg as a variable.
 To achieve this, we explore the relationship between the multi-view rendered images.
-A random image is selected from all rendered images as the reference image, and the cosine similarity between the other images and the reference image is computed.
+First, each rendered image is encoded into high-dimensional semantic features, which capture the global view-aware semantics of the images and exhibit some robustness to flipping, rotation, and symmetry variations of the rendered content [bengio2014representationlearningreviewnew].
+A random image is selected from all rendered images as the reference image, and the cosine similarity between the other images and the reference image is computed at the feature level.
 As shown in Fig.6, the red camera represents the reference view selected from a series of randomly generated camera views $\Lambda$.
@@
 The reference camera is then mirrored to a symmetric point using this line.
-The azimuthal distances between other camera views and the nearest reference camera view are computed, determining the expected similarity partial order among the corresponding rendered images: $sim_{0,3}>sim_{0,2}>sim_{0,1}$.
-The similarity relationship between the 2D projections obtained from Eq.(8) must also adhere to this order.
-Accordingly, the cosine similarity between the corresponding rendered images is computed as the actual similarity and a similarity-based partial order loss $\mathcal{L}_{P}$ is employed to constrain the actual similarity to align with the expected similarity relationships.
-$\mathcal{L}_{P}$ is defined as:
-$\mathcal{L}_{P}\!=\!\sum_{i=1}^{|\Lambda|-2}\!\max\!\left(0,{sim}\left(R^{i+1},R^{0}\right)\!-\!{sim}\left(R^{i},R^{0}\right)\right),$ (16)
-where $R^{i}$ is the 2D rendered image from the 3D content with the $i^{th}$ view in $\Lambda$, $R^{0}$ is the reference image and $\operatorname{sim}\left(R^{i},R^{0}\right)$ denotes the cosine similarity between $R^{i}$ and $R^{0}$.
-The $\mathcal{L}_{P}$ term can constrain the multi-face Janus problem early in training.
-As shown in Fig.6, using “A squirrel eating a burger” as an example, without the $\mathcal{L}_{P}$ constraint, the generated result exhibits an irreconcilable multi-face Janus problem.
-When $\mathcal{L}_{P}$ is applied, the generated 3D content shows significant improvement starting from the 1000th iteration.
-Subsequently, we utilize a parameter update module based on DDIM and ISM, which originates from the methodologies outlined in LucidDreamer.
-Specifically, with a given prompt $c$ and the noisy latents $\boldsymbol{z}_{s}$ and $\boldsymbol{z}_{t}$ generated through DDIM inversion from $\boldsymbol{z}_{0}$ with the time interval $t-s$, the ISM loss is defined as:
-$\mathcal{L}_{ISM}(\theta)\!=\!\mathbb{E}_{t,c}\!\left[\omega(t)\!\left\|\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{z}_{t},t,c\right)\!-\!\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{z}_{s},s,\emptyset\right)\right\|_{2}^{2}\right].$ (17)
-For detailed information of ISM, please refer to [17]. The final loss function is then formalized as:
-${Loss}=\mathcal{L}_{ISM}+\kappa\mathcal{L}_{P},$ (18)
-where $\kappa$ is the weight coefficients of $\mathcal{L}_{P}$.
+The azimuthal distances between other camera views and the nearest reference camera view are computed, determining the expected similarity partial order among the corresponding rendered images: $\text{sim}_{0,3}>\text{sim}_{0,2}>\text{sim}_{0,1}$.
+The similarity relationship between the features of rendered images obtained from Eq.(9) must also conform to this order.
+Accordingly, the cosine similarity between the semantic features of corresponding rendered images is computed as the actual similarity and a similarity-based partial order loss $\mathcal{L}_{\text{P}}$ is employed to constrain the actual similarity to align with the expected similarity relationships.
+$\mathcal{L}_{\text{P}}$ is defined as:
+$\mathcal{L}_{\text{P}}\!=\!\sum_{i=1}^{|\Lambda|-2}\!\max\!\left(0,\text{sim}\!\left(F^{i+1},F^{0}\right)\!-\!\text{sim}\!\left(F^{i},F^{0}\right)\right),$ (17)
+TABLE II: Quantitative Results of Ablation Study
+[…]
+[…]
+Rank Score Rank Score Rank Score Rank Score Rank Score Rank Score
+Janus Issue 7 -0.770 1 0.340 7 0.086 7 0.050 7 100.0 7 100.0
+Baseline 6 -0.637 2 0.338 6 0.079 6 0.047 6 85.7 6 97.1
+Baseline + $\mathcal{L}_{\text{P}}$ 5 0.274 7 0.327 3 0.074 3 0.042 5 37.1 4 60.0
+Baseline + Perp-Neg 4 0.285 6 0.328 4 0.076 5 0.043 4 24.0 5 64.0
+Baseline + Perp-Neg + $\mathcal{L}_{\text{P}}$ 3 0.294 5 0.332 4 0.076 3 0.042 3 15.3 3 52.7
+Baseline + VDM 2 0.465 3 0.333 2 0.068 2 0.039 2 11.4 2 42.0
+Baseline + VDM + $\mathcal{L}_{\text{P}}$ (Ours) 1 0.707 3 0.333 1 0.060 1 0.035 1 3.4 1 31.0
+where $F^{i}$ is the encoded feature of the rendered image corresponding to the $i^{th}$ view in $\Lambda$, $F^{0}$ is the reference feature and $\operatorname{sim}\left(F^{i},F^{0}\right)$ denotes the cosine similarity between $F^{i}$ and $F^{0}$.
+The $\mathcal{L}_{\text{P}}$ term can constrain the multi-face Janus problem early in training.
+As shown in Fig.7, using “A squirrel eating a burger” as an example, without the $\mathcal{L}_{\text{P}}$ constraint, the generated result exhibits an irreconcilable multi-face Janus problem.
+When $\mathcal{L}_{\text{P}}$ is applied, the generated 3D content shows significant improvement starting from the 1000th iteration.
+Our plug-and-play method integrates into multiple baseline frameworks with diverse score distillation strategies, unifying their score distillation losses into a consistent formulation $\mathcal{L}_{\text{score}}$.
+The final loss function is then formalised as:
+${Loss}=\mathcal{L}_{\text{score}}+\kappa\mathcal{L}_{\text{P}},$ (18)
+where $\kappa$ is the weight coefficient of $\mathcal{L}_{\text{P}}$.
 IV Experiments
@@
 First, we conduct a comprehensive ablation study in Section IV-A to evaluate the effectiveness of each individual module.
-Next, Section IV-B presents comprehensive comparisons against seven SOTA methods to further highlight the advantages of our framework.
-Finally, we explore the improvement brought by VDM to T2I tasks in Section IV-C, further extending the evaluation.
-All experiments were conducted using a single NVIDIA RTX 4090 GPU and the Stable Diffusion v2.1
-Figure 7: Qualitative results of ablation study.
-The first row shows the results without the intervention of VDM or $\mathcal{L}_{P}$.
-The second and third rows display the generated results after applying VDM or $\mathcal{L}_{P}$ respectively.
-The bottom row presents the performance with both modules jointly optimized.
-Each generated object is examined from different angles to visually assess its multi-view consistency.
-base
+Furthermore, Perp-Neg is additionally considered as an ablation variable to ensure a fair and comprehensive comparison.
+Next, in Section IV-B, ConsDreamer is integrated into six SOTA baselines encompassing diverse 3D representations (NeRF or 3DGS) and various score distillation strategies to validate the versatility of our method.
+Finally, we explore the improvement brought by VDM to T2I tasks in Section IV-D, further extending the evaluation.
+All experiments were conducted using the Stable Diffusion v2.1 base model for distillation.
+Figure 9: Ablation on weight coefficient $\kappa$ of $\mathcal{L}_{\text{P}}$.
+$\kappa$ = 0.6 yields the highest visual quality: excessively large values lead to over-smoothing and misalignment with textual prompts (e.g., “white hair”), while overly small values cause view inconsistency in the generated results, resulting in poor fitting of the target object.
+IV-A Ablation Studies
+1) Qualitative Results:
 Figure 10: Examples of inconsistencies in generated 3D content.
@@
 Additionally, the section below showcases other types of inconsistencies, including component redundancy and component morphing.
-model for distillation. Besides, we set the training process with 5,000 iterations and employed the pre-trained Point-E [31] to initialize the 3D Gaussian Splatting.
-IV-A Ablation Studies
-1) Qualitative Results:
-TABLE I: Quantitative Results of Ablation Study on ConsDreamer
-[…]
-[…]
-Rank Score Rank Score Rank Score Rank Score Rank Rate Rank Rate
-Janus issue 5 -0.770 1 0.340 5 0.086 5 0.050 5 100.0 5 100.0
-w/o $\mathcal{L}_{p}$ w/o VDM 4 0.282 5 0.328 3 0.076 4 0.043 4 24.0 4 64.0
-w/o VDM 3 0.294 4 0.332 3 0.076 3 0.042 3 15.3 3 52.7
-w/o $\mathcal{L}_{p}$ 2 0.465 2 0.333 2 0.068 2 0.039 2 11.4 2 42.0
-Ours 1 0.707 2 0.333 1 0.060 1 0.035 1 3.4 1 31.0
-  • $\uparrow$
-    Higher scores indicate better performance.
-  • $\downarrow$
-    Lower scores indicate better performance.
 By eliminating view semantic bias in T2I mappings and incorporating explicit consistency constraints across multi-view images, our method significantly improves the consistency of generated content.
-As shown in Fig.7, both VDM and the perceptual loss $\mathcal{L}_{P}$ effectively reduce the occurrence of the multi-face Janus issue.
-However, certain inconsistency artifacts persist, as demonstrated in the second and third rows of Fig.7, including redundant Batman ears, additional Pikachu limbs, and extraneous nuclear reactors on Iron Man’s back.
-Notably, when these two modules are combined, the proposed method achieves highly consistent 3D content, surpassing the performance of using either module individually.
-The results in the bottom row of Fig.7 demonstrate that our approach not only mitigates the multi-face Janus issue but also effectively prevents component redundancy.
+As illustrated in Fig.8, the baseline model without any consistency strategy suffers from severe multi-face Janus artefacts, which are characterised by additional limbs and redundant faces of Pikachu.
+When the perceptual loss $\mathcal{L}_{\text{P}}$ is applied, the extra limbs and ears are effectively suppressed.
+To investigate the performance gain of VDM over Perp-Neg, we also treat Perp-Neg as an ablation variable, and the visualisation results demonstrate a notable improvement with VDM achieving more consistent generation results than Perp-Neg.
+However, certain inconsistent artefacts, such as redundant limbs still persist.
+The model yields the optimal performance in terms of both consistency and visual quality when VDM is combined with $\mathcal{L}_{\text{P}}$.
+Furthermore, as shown in Fig.
+9, we conducted an ablation study on the weight coefficient $\kappa$ of the partial order loss $\mathcal{L}_{\text{P}}$.
 Figure 11: Similarity distribution between prompts and rendered results across methods.
 The similarity distribution between prompts and rendered results for various methods is compared, the results demonstrate that our method generates 3D content with a similarity score distribution that aligns more closely with the ideal distribution.
-Figure 10: Qualitative comparisons of 3D content generated by different methods.
-This figure presents a comparison of our method with seven SOTA text-to-3D generation approaches across various objects, evaluated from five viewpoints (0°, 45°, 90°, 135°, and 180°) for quality and consistency.
-When generating Super Saiyan characters, DreamFusion, Sherpa3D, LucidDreamer, ScaleDreamer, DreamScene, and GCS-BEG produced redundant limbs.
-For Iron Man, Sherpa3D, LucidDreamer, ScaleDreamer, and DreamScene incorrectly generated a nuclear reactor on the back.
-In the case of Batman, Sherpa3D, LucidDreamer, DreamScene, and GCS-BEG generated three ears.
-The results demonstrate that our proposed ConsDreamer outperforms SOTA methods in terms of both generation quality and multi-view consistency.
-TABLE II: Quantitative Comparison with SOTA Methods
-[…]
-[…]
-Rank Score Rank Score Rank Score Rank Score Rank Score Rank Rate Rank Rate
-DreamFusion [14] 8 3.48 8 4.37 8 3.93 7 -0.289 6 0.307 8 60.0 8 86.7
-Debiased-SDS [15] 7 4.19 7 4.93 7 4.56 8 -0.563 8 0.282 4 46.7 2 60.0
-Sherpa3D [22] 6 5.42 6 5.82 6 5.62 6 -0.161 7 0.291 7 56.3 7 75.0
-LucidDreamer [17] 4 6.76 5 6.31 4 6.54 4 0.285 4 0.328 2 24.0 3 64.0
-ScaleDreamer [24] 2 7.82 4 6.33 3 7.08 2 0.371 1 0.333 5 47.3 6 69.2
-DreamScene [23] 5 6.25 3 6.90 5 6.58 5 0.201 5 0.321 6 48.3 5 66.7
-GCS-BEG [21] 3 7.03 2 7.65 2 7.34 3 0.333 3 0.316 3 33.3 4 64.3
-Ours 1 7.85 1 8.05 1 7.95 1 0.707 1 0.333 1 3.4 1 31.0
-  • $\uparrow$
-    Higher scores indicate better performance.
-  • $\downarrow$
-    Lower scores indicate better performance.
-2) Quantitative Results: Qualitative results visually demonstrate the effectiveness of the proposed framework.
-We employed conventional evaluation metrics in text-to-3D generation, including ImageReward [50], OpenCLIP-L/14 [51], and A-LPIPS [52], to objectively assess the effectiveness of our framework.
+2) Quantitative Results: Qualitative results visually demonstrate the effectiveness of the proposed method.
+We employed conventional evaluation metrics in text-to-3D generation, including ImageReward [NEURIPS2023_33646ef0-imagereward], OpenCLIP-L/14 [ramesh2022hierarchical-openclip], and A-LPIPS [Zhang_2018_CVPR-alpips], to objectively assess the effectiveness of our method.
 ImageReward measures the consistency between generated results and text descriptions, OpenCLIP-L/14 provides enhanced text-image alignment capabilities to capture fine-grained semantic information, and A-LPIPS evaluates the visual quality of generated images, focusing on perceptual similarity.
-Table I shows that each module of our method outperforms the scenario without intervention in various aspects, and the joint optimization of both modules achieves even better performance.
-Additionally, we evaluated the percentage of generated results exhibiting Inconsistency, denoted as $f_{inc}$.
-We broadened the traditional definition of the multi-face Janus problem to encompass a wider range of inconsistencies, such as extra limbs and other implausible artifacts.
+Table II shows that each module of our method outperforms the scenario without intervention in various aspects, and the joint optimisation of both modules achieves even better performance.
+Additionally, we evaluated the percentage of generated results exhibiting Inconsistency, denoted as $\text{f}_{\text{inc}}$.
+We broadened the traditional definition of the multi-face Janus problem to encompass a wider range of inconsistencies, such as extra limbs and other implausible artefacts.
 Fig.
@@
 For a comprehensive evaluation, we selected a diverse set of prompts involving objects with distinct front-back differences (e.g., portraits, animals, and vehicles).
-The quantitative results are summarized in Table I, where the frequency of the multi-face Janus problem is denoted as $f_{mf}$.
-The results indicate that any of our proposed components significantly improves the consistency of the generated content.
-Specifically, applying $\mathcal{L}_{p}$ achieves significant improvements, reducing the multi-face problem and enhancing overall consistency by 36.25% and 18.75%, respectively, compared to the scenario without intervention.
-Similarly, applying VDM demonstrates further enhancements, with improvements of 52.5% and 34.38%, respectively.
-When both modules are jointly applied, the performance gains are substantially amplified, reaching 85.83% and 51.56%, respectively.
-Figure 11: Application of the VDM to T2I Tasks.
-The results demonstrate the outputs of Stable Diffusion with explicit view descriptions across six prompts.
-The first column in each set shows the original outputs, highlighting the inherent view bias in T2I models.
-By applying the VDM, the view features were disentangled, enabling precise control of desired views (e.g., “back view” or “side view”).
-However, we chose those generation results with the multi-face Janus problem and observed that the OpenCLIP-L/14 metric yields higher scores in Table I.
+The quantitative results are summarised in Table II, where the frequency of the multi-face Janus problem is denoted as $\text{f}_{\text{mf}}$.
+The results indicate that each proposed component significantly improves the consistency of the generated content.
+Notably, when directly comparing VDM against Perp-Neg, VDM demonstrates substantially superior performance, achieving an additional 52.5% reduction in $\text{f}_{\text{mf}}$ (from 24.0% to 11.4%) and a 34.4% reduction in $\text{f}_{\text{inc}}$ (from 64.0% to 42.0%).
+This significant improvement validates VDM’s effectiveness in addressing prior view biases.
+When combined with $\mathcal{L}_{\text{P}}$, the performance gains are further amplified.
+Our complete method (VDM + $\mathcal{L}_{\text{P}}$) achieves the best overall performance with $\text{f}_{\text{mf}}$ of only 3.4% and $\text{f}_{\text{inc}}$ of 31.0%, representing reductions of 96.0% and 68.1% compared to the baseline, respectively.
+Compared to Perp-Neg + $\mathcal{L}_{\text{P}}$, our method further reduces $\text{f}_{\text{mf}}$ by 77.8% and $\text{f}_{\text{inc}}$ by 41.2%, demonstrating the synergistic effects of VDM and $\mathcal{L}_{\text{P}}$ in effectively mitigating the multi-face Janus problem.
+Figure 12: Qualitative text-to-3D generation results of our method.
+We integrate ConsDreamer into six SOTA baselines spanning NeRF/3DGS representations and diverse score distillation strategies, and these results demonstrate that our proposed method is visually superior to the baselines.
+However, we chose those generation results with the multi-face Janus problem and observed that the OpenCLIP-L/14 metric yields higher scores in Table II.
 Fig.11 presents the OpenCLIP similarity distribution between prompts and rendered results for different methods, including samples exhibiting the multi-face Janus problem for comparison.
@@
 The results demonstrate that our method produces a similarity score distribution that aligns more closely with the ideal one.
-IV-B Comparison with SOTA Methods
-1) Qualitative Results: Fig.2 showcases the vivid and diverse text-to-3D results generated by our method.
-Subsequently, a qualitative comparison is conducted against seven SOTA text-to-3D generation methods.
-To better assess the quality and view consistency of the generated models, five evenly distributed viewpoints over a 180° azimuth rotation are evaluated.
+TABLE III: Quantitative Results of Integrating ConsDreamer into Various Baselines
+Methods 3D Representation Score Distillation Consistency Strategy User Study ImageReward $\uparrow$ OpenCLIP $\uparrow$ Frequency of Inconsistency Time (min) VRAM (GiB)
+[…]
+DreamFusion NeRF SDS \ 3.48 4.37 3.93 $0.289\pm 0.81$ $0.307\pm 0.025$ 60.0 86.7 $\sim$ 40 6.2
++ Ours NeRF SDS ConsDreamer 5.45 5.58 5.52 $\textbf{0.103}\pm 0.79$ $\textbf{0.317}\pm 0.023$ 26.7 60.0 $\sim$ 42 6.3
+SDI NeRF SDI Perp-Neg 6.33 6.89 6.61 $0.297\pm 0.61$ $0.328\pm 0.024$ 33.3 61.9 $\sim$ 119 39.2
++ Ours NeRF SDI ConsDreamer 6.51 7.03 6.77 $\textbf{0.175}\pm 0.59$ $\mathbf{0.331}\pm 0.021$ 27.0 40.5 $\sim$ 116 39.0
+DreamScene 3DGS FPS Perp-Neg 6.25 6.90 6.58 $0.201\pm 0.75$ $0.321\pm 0.027$ 48.3 66.7 $\sim$ 30 13.6
++ Ours 3DGS FPS ConsDreamer 6.93 7.45 7.19 $\mathbf{0.531}\pm 0.52$ $\textbf{0.324}\pm 0.027$ 28.3 45.0 $\sim$ 30 13.6
+LucidDreamer 3DGS ISM Perp-Neg 6.76 6.31 6.54 $0.285\pm 0.69$ $0.328\pm 0.019$ 24.0 64.0 $\sim$ 55 22.1
++ Ours 3DGS ISM ConsDreamer 7.85 8.05 7.95 $\mathbf{0.707}\pm 0.41$ $\mathbf{0.333}\pm 0.015$ 3.4 31.0 $\sim$ 55 22.0
+GCS-BEG 3DGS GCS Perp-Neg 7.03 7.65 7.34 $0.333\pm 0.83$ $0.316\pm 0.031$ 33.3 64.3 $\sim$ 41 16.4
++ Ours 3DGS GCS ConsDreamer 7.97 8.16 8.07 $\mathbf{0.814}\pm 0.51$ $\textbf{0.327}\pm 0.035$ 7.1 35.7 $\sim$ 41 16.3
+GE3D 3DGS GE3D Perp-Neg 6.58 6.93 6.76 $0.297\pm 0.47$ $0.324\pm 0.020$ 34.3 65.7 $\sim$ 109 12.5
++ Ours 3DGS GE3D ConsDreamer 7.01 7.64 7.33 $\mathbf{0.447}\pm 0.44$ $\textbf{0.324}\pm 0.019$ 17.1 36.7 $\sim$ 109 12.5
+IV-B Integration with SOTA Methods
+1) Qualitative Results: Fig.2 presents vivid and diverse text-to-3D generation results produced by our method.
+Furthermore, we integrate ConsDreamer into a range of advanced 3D generation frameworks, which cover different 3D representation paradigms and various score distillation strategies.
 As demonstrated in Fig.12, our method substantially alleviates the multi-face Janus issue prevalent in existing approaches.
-2) Quantitative Results: We selected ImageReward, OpenCLIP-L/14, and the frequency of Inconsistency as objective evaluation metrics.
-As shown in Table II, our method achieves the highest scores across all metrics, particularly in the rate of the multi-face Janus issue, where our method outperforms the average by approximately 90%.
-We conducted a user study involving 105 participants to compare with other SOTA methods as subjective evaluation results.
-Using a 1 to 10 scale, participants assessed the models based on two key criteria: “Generation Quality” and “ 3D Consistency.” To ensure a fair and unbiased comparison, all samples used in the study were randomly selected, with no cherry-picking to favor our approach.
-As shown in Table II, our method achieves the highest average scores across both evaluation metrics compared to the seven competing methods.
+Figure 13: Application of the VDM to T2I Tasks.
+The results demonstrate the outputs of Stable Diffusion with explicit view descriptions across six prompts.
+The first column in each set shows the original outputs, highlighting the inherent view bias in T2I models.
+By applying the VDM, the view features were disentangled, enabling precise control of desired views (e.g., “back view” or “side view”).
+2) Quantitative Results: We selected ImageReward, OpenCLIP-L/14, and inconsistency frequency as objective evaluation metrics.
+As shown in Table III, ConsDreamer delivers more substantial gains over the baseline than PerpNeg, attaining state-of-the-art performance across all metrics.
+Specifically, in the multi-face Janus issue rate, our method outperforms PerpNeg-based baselines by approximately 52.3%.
+Additionally, we report the mean runtime and GPU VRAM usage.
+For GPU VRAM usage, we compute the average of the maximum GPU memory consumption across multiple runs.
+Table III demonstrates that our method is lightweight, as it does not compromise GPU memory or runtime for consistent generation.
+We conducted a user study involving 145 participants to compare with other SOTA methods as subjective evaluation results.
+Using a 1 to 10 scale, participants assessed the models based on two key criteria: “Generation Quality” and “ 3D Consistency.” To ensure a fair and unbiased comparison, all samples used in the study were randomly selected, with no cherry-picking to favour our approach.
+As shown in Table III, our method achieves the highest average scores across both evaluation metrics compared to baselines.
 These results demonstrate that the 3D content generated by our approach is consistently more appealing and exhibits superior view consistency.
@@
 The results clearly demonstrate that the VDM mitigates view semantic errors and significantly improves generation accuracy.
-2) Quantitative Results: We compared the success rate of generating images with specific views using Stable Diffusion and our method.
+2) Quantitative Results: We compared the success rate of generating images with specific views using Stable Diffusion, Perp-Neg and our method.
 Given the open-ended nature of the generated content, a strict definition was adopted for counting success cases.
-Images exhibiting hallucinations (e.g., artifacts or incorrect semantics) were excluded.
+Images exhibiting hallucinations (e.g., artefacts or incorrect semantics) were excluded.
 Additionally, in instances where multiple entities appeared in an image, it was considered a failure if even one entity did not match the specified view.
-The results, presented in Table III, demonstrate that our method achieves a significantly higher success rate compared to Stable Diffusion, underscoring its effectiveness in aligning multimodal view semantics.
-TABLE III: Success Rate of Specific View Generation
-Methods Successful Generation Rate(%$\uparrow$
+The results, presented in Table IV, show that our proposed VDM improves rear-view generation success by approximately 36.7% over Perp-Neg, demonstrating that our method achieves a substantially higher success rate and effectively aligns multimodal view semantics.
+Figure 14: Generation failure case examples.
+For complex or multi-object prompts (top), the underlying optimisation may still produce distorted or entangled geometry, while for single-object cases (middle and bottom), ConsDreamer improves view consistency but cannot fully remove illumination artefacts such as oversaturation and colour shifts.
+TABLE IV: Success Rate of Specific View Generation
+[…]
 Side view Back view
-Stable Diffusion [34] 66.3 28.7
-Stable Diffusion+VDM 80.2 77.2
-  • $\uparrow$
-    Higher scores indicate better performance.
+Stable Diffusion 66.3 28.7
+Stable Diffusion + Perp-Neg 77.4 56.5
+Stable Diffusion + VDM (Ours) 80.2 77.2
+IV-D Limitations
+Methods built on score-distillation pipelines, while capable of substantially improving multi-view consistency, still inherit the fundamental limitations of this optimisation paradigm.
+For highly complex text descriptions or cluttered multi-object scenes, the underlying optimisation may converge to sub-optimal 3D configurations, leading to fragmented or entangled structures.
+In these regimes, partial-order constraints act only as a relatively weak regularisation signal and cannot by themselves guarantee globally coherent geometry, as illustrated in Fig.
+14.
+Moreover, score-distillation-based approaches predominantly target view-level semantics and geometry, and typically do not explicitly account for illumination artefacts introduced during distillation.
+Brightness accumulation and colour saturation anomalies can therefore persist even when view consistency is improved (Fig.
+14).
+A principled treatment of such photometric effects would likely require revisiting the distillation objective itself, which we regard as an important direction for future work.
 V Conclusion
 In this paper, we conducted a mathematical analysis of the multi-face Janus problem in zero-shot text-to-3D generation.
-To address this challenge, we proposed a novel framework, $ConsDreamer$, which incorporates a View Disentanglement Module to effectively eliminate prior view biases and enhance view control.
-Additionally, we designed a Partial Order Loss $\mathcal{L}_{P}$ to explicitly enforce multi-view consistency across generated content.
-We validated the significant improvements of our method over SOTA approaches through comprehensive qualitative and quantitative evaluations.
-Beyond its application in 3D generation, the View Disentanglement Module also proves effective in 2D generation tasks, correcting view-related semantic biases during the image synthesis process.
-These contributions collectively establish a robust framework for generating 3D content with greater consistency and realism, paving the way for broader applications in creative industries, virtual reality, and beyond.
+While Previous methods alleviate this problem by orthogonalizing negative prompts in the denoising score space, they only operate on the UNet output and overlook the aggregated prior view features formed during the UNet processing, leading to the recurrence of inconsistency when lifting 2D representations to 3D.
+To address this fundamental limitation, we propose ConsDreamer, which offers two key advantages: (1) a View Disentanglement Module that eliminates view biases in conditional prompts by decoupling irrelevant view components and injecting precise view control prior to denoising; and (2) a similarity-based partial order loss that enforces geometric consistency in the unconditional term by aligning cosine similarities with azimuthal relationships, synergistically enhancing view semantic clarity.
+Demonstrating strong versatility, ConsDreamer can be seamlessly integrated into diverse 3D representations (e.g., NeRF, 3DGS) and various score distillation paradigms.
+Extensive experiments across multiple SOTA baselines confirm that our method effectively mitigates the multi-face Janus problem while outperforming existing approaches in both visual quality and multi-view consistency.